### v1

# Active Human Feedback Collection via Neural Contextual Dueling Bandits

Abstract
Collecting human preference feedback is often expensive, leading recent works to develop principled algorithms to select them more efficiently. However, these works assume that the underlying reward function is linear, an assumption that does not hold in many real-life applications, such as online recommendation and LLM alignment. To address this limitation, we propose Neural-ADB, an algorithm based on the neural contextual dueling bandit framework that provides a principled and practical method for collecting human preference feedback when the underlying latent reward function is non-linear. We theoretically show that when preference feedback follows the Bradley-Terry-Luce model, the worst sub-optimality gap of the policy learned by Neural-ADB decreases at a sub-linear rate as the preference dataset increases. Our experimental results on problem instances derived from synthetic preference datasets further validate the effectiveness of Neural-ADB.
1 Introduction
Collecting human preference feedback is essential in many real-life applications, like online recommendations (Kohli et al., 2013; Wu et al., 2023; Zhang & Wang, 2023; Yang et al., 2024), content moderation (Avadhanula et al., 2022), medical treatment design (Lai & Robbins, 1985; Bengs et al., 2021), prompt optimization (Lin et al., 2024), and aligning large language models (Bai et al., 2022; Mehta et al., 2023; Menick et al., 2022; Chaudhari et al., 2024; Das et al., 2024; Ji et al., 2024), to ensure systems effectively align with user preferences and exhibit desired behaviors. However, this process is often costly due to the need for skilled evaluators, the complexity of tasks, and the time-intensive nature of producing high-quality, reliable human feedback. To address the challenge of balancing cost and effectiveness in aligning systems, this paper proposes principled and practical algorithms for efficiently collecting human feedback sequentially and adaptively to achieve the desired system behavior. Specifically, this paper aims to answer the following fundamental question:
How to achieve desired system behavior while using as minimum human feedback as possible?
Recent works (Mehta et al., 2023; Das et al., 2024) have modeled the problem of active human feedback collection as an active version of the contextual dueling bandit problem (ADB for brevity) (Saha, 2021; Bengs et al., 2022; Li et al., 2024), where context-arm pair in the contextual dueling bandits corresponds to a task for which human preference feedback is collected and then proposed algorithms to select context-arm pairs for human feedback sequentially and adaptively by exploiting collected preference dataset, i.e., past context-arm pairs with their preference feedback. The preference feedback between two context-arm pairs is commonly assumed to follow the Bradley-Terry-Luce (BTL) model [^1] For more than two context-arm pairs, preferences are typically modeled using the Plackett-Luce model (Soufiani et al., 2014). (Hunter, 2004; Luce, 2005; Saha, 2021; Bengs et al., 2022; Li et al., 2024; Lin et al., 2024; Verma et al., 2025) in which the probability of preferring a context-arm pair over others is proportional to the exponential of its reward. In many real-life applications, the number of context-arm pairs (e.g., prompt-response pair in prompt optimization or user-movie pair in online movie recommendation) can be large or even infinite. Therefore, the reward for each context-arm pair is assumed to be parameterized by an unknown function of its feature vector.
To better align the system for optimal performance, we consider two key components: context selection and arm selection. The context selection aims to encourage diversity by exploring the context space, such as selecting prompts as diverse as possible in prompt optimization. Whereas arm selection focuses on identifying the arms that help learn the best arm for each context, such as selecting the most effective pair of responses to a given prompt that maximizes the system’s learning (Lin et al., 2024; Verma et al., 2025). Since the goal is to identify the best arm for each context, selecting suboptimal arms provides less useful information than choosing better arms. Existing methods for active contextual dueling bandits (Mehta et al., 2023; Das et al., 2024) fail to incorporate an efficient arm selection strategy during the data collection process, thereby limiting the ability of these methods to achieve optimal performance.
An efficient arm selection strategy requires estimating the reward function to guide the arm selection process effectively. Since the reward function may not always be linear in practice, this paper parameterizes the reward function via a non-linear function, which needs to be estimated using the available preference dataset. This non-linear function can be estimated by using methods like Gaussian processes (Williams & Rasmussen, 2006; Srinivas et al., 2010) or neural networks (Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2025). However, Gaussian processes have limited expressive power and fail to optimize highly complex functions. In contrast, neural networks (NNs) have greater expressive power, making them well-suited for modeling complex functions (Dai et al., 2023; Lin et al., 2023; 2024; Verma et al., 2025).
In this paper, we propose a neural active contextual dueling bandit algorithm, Neural-ADB, which uses an NN to estimate the unknown reward function using the available preference dataset. The context selection in Neural-ADB is adapted from Das et al. (2024), while arm selection strategies are based on, respectively, upper confidence bound (UCB) and Thompson sampling (TS), and adapted from Verma et al. (2025). Due to the differences in context selection strategy, arm selection strategies, and the use of a non-linear reward function, our theoretical analysis is completely different than related existing work (Mehta et al., 2023; Das et al., 2024). One of the key theoretical contributions of this paper is providing an upper bound on the maximum Mahalanobis norm of a vector from the fixed input space, measured with respect to the inverse of a positive definite Gram matrix that is constructed using finite, adapted samples from that space. Building on this result, we prove that the worst sub-optimality gap (defined in Eq. 1) of the policy learned by Neural-ADB decreases at a sub-linear rate as the preference dataset size increases. Specifically, our contribution are as follows:
  • We introduce the setting of active contextual dueling bandits with a non-linear reward function in Section 2. In Section 3, we propose a practical neural active contextual dueling bandit algorithm, Neural-ADB, which uses an NN to estimate the unknown reward function from the available preference dataset and then uses this estimate into the arm selection strategies.
  • One of the main theoretical contributions of this paper is proving an upper bound on the maximum Mahalanobis norm of a vector from the fixed input space, as measured with respect to the inverse of a positive definite Gram matrix (1), where the gram matrix is constructed using finite, adapted samples from that input space. We show that this upper bound decays at a sub-linear rate as the number of samples used in the Gram matrix increases. This result itself is of independent interest, as it gives valuable insights beyond the specific application of our work.
  • We prove that the worst sub-optimality gap of the policy learned by Neural-ADB with both of our arm selection strategies (2 and 3) decreases at a sub-linear rate with respect to the size of preference dataset, specifically at rate of $\tilde{O}((\tilde{d}/T)^{\frac{1}{2}})$, where $\tilde{O}$ hides the logarithmic factors and constants, and $\tilde{d}$ is the effective dimension of context-arm feature vectors. The decay rate of the worst sub-optimality gap for Neural-ADB improves by a factor of $\tilde{O}(({\tilde{d}\log T})^{\frac{1}{2}})$ compared to exiting algorithms (Mehta et al., 2023; Das et al., 2024), thus bridging the gap between theory and practice.
  • Finally, in Section 4, our experimental results further validate the different performance aspects of Neural-ADB.
2 Problem Setting
We formulate active human preference feedback collection as an active contextual dueling bandit problem, where a labeler (human or simulator) provides preference feedback for a chosen pair of options (arms), such as prompt-response pairs in prompt optimization (Lin et al., 2024) and LLM alignment (Mehta et al., 2023; Das et al., 2024).
Active contextual dueling bandit problem.   We consider an active contextual dueling bandit problem, where the underlying latent reward function can be non-linear. In each iteration of this problem, the learner’s goal is to select a triplet containing a context and two arms for collecting preference feedback from a labeler/human such that the collected preference dataset leads to superior performance. Let $\mathcal{C}$ be the set of contexts and $\mathcal{A}$ be the set of all possible arms. In each iteration, the learner selects a context $c_{t}\in\mathcal{X}$ and then two arms (denoted as $a_{t,1}$ and $a_{t,2}$) from the set of arms $\mathcal{A}$. After selecting the triplet of context and two arms, the learner receives a stochastic preference feedback $y_{t}$, where $y_{t}=1$ implies the arm $a_{t,1}$ is preferred over arm $a_{t,2}$ for the context $c_{t}$ and $y_{t}=0$ otherwise. We use $\varphi(c_{t},a)$ to denote the context-arm feature vector for context $c_{t}$ and an arm $a$, where $\varphi:\mathcal{C}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ is a known feature map.
Preference model.   Following the dueling bandits literature (Saha, 2021; Bengs et al., 2022; Li et al., 2024; Verma et al., 2025), we assume the preference feedback follows the Bradley-Terry-Luce (BTL) model [^2] Our results are also applicable to any preference models, such as the Thurstone-Mosteller model and Exponential Noise, as long as stochastic transitivity holds (Bengs et al., 2022). (Hunter, 2004; Luce, 2005). Under the BTL preference model, the preference feedback has a Bernoulli distribution, where the probability that the first selected arm $a_{t,1}$ is preferred over the second selected arm $a_{t,2}$ for the given context $c_{t}$ is given by
$\mathbb{P}\left\{a_{t,1}\succ a_{t,2}\right\}$ =$=\mathbb{P}\left\{y_{t}=1|c_{t},a_{t,1},a_{t,2}\right\}=\mu\left(f(\varphi(c_{t},a_{t,1}))-f(\varphi(c_{t},a_{t,2}))\right),$
where $a_{t,1}\succ a_{t,2}$ used for brevity and denotes that $a_{t,1}$ is preferred over $a_{t,2}$ for the given context $c_{t}$, $\mu(x)=1/(1+\mathrm{e}^{-x})$ is the sigmoid function, $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is an unknown non-linear bounded reward function, and $f(\varphi(c,a))$ is the latent reward of the arm $a$ for the context $c$. We require the following standard assumptions on the function $\mu$ (commonly referred to as a link function in the bandit literature (Li et al., 2017; Bengs et al., 2022)):
Assumption 1.
  • Let $\kappa_{\mu}\doteq\inf\limits_{c\in\mathcal{C},a,b\in\mathcal{A}}\dot{\mu}(f(\varphi(c,a))-f(\varphi(c,b)))>0$ for all triplets of context and arms.
  • The link function $\mu:\mathbb{R}\rightarrow[0,1]$ is continuously differentiable and Lipschitz with constant $L_{\mu}$. For logistic function, we have $L_{\mu}\leq 1/4$.
Performance measure.  We denote the collected preference dataset up to $T$ iterations by $\mathcal{D}_{T}=\{(c_{s},a_{s,w},a_{s,l},y_{s})\}_{s=1}^{T}$, where $a_{s,w}\succ a_{s,l}$ for the selected context $c_{s}$ in iteration $s$. We aim to learn a policy, $\pi:\mathcal{C}\rightarrow\mathcal{A}$ from the collected preference dataset $\mathcal{D}_{T}$ that achieves the worst suboptimality across all contexts in $\mathcal{C}$, which is defined as follows:
$\Delta_{\mathcal{D}_{T}}^{\pi}=\max_{c\in\mathcal{C}}\left[\max_{a\in\mathcal{A}}f(\varphi(c,a))-f\left(\varphi(c,\pi(c))\right)\right],$ (1)
where policy $\pi$ is a learned policy from the collected preference dataset $\mathcal{D}_{T}$ up to the iteration $T$. The policy $\pi_{\mathcal{D}_{T}}$ competes with the Condorcet winner (Bengs et al., 2021; Das et al., 2024) for a given context, i.e., an arm that is better than all other arms. The suboptimality gap is the worst possible difference in latent rewards over the set of contexts, and the same performance measure is used in prior work (Mehta et al., 2023; Das et al., 2024).
3 Algorithm for Active Human Preference Feedback Collection
This section introduces Neural-ADB, a simple yet principled and practical algorithm designed to efficiently select context-arm pairs for collecting preference feedback. Neural-ADB consists of two main components: context selection and arm selection. Since the arm selection strategy depends on the estimated reward function, we first explain how an NN can be used to estimate the unknown reward function. We will then give details of the context and arm selection strategies, followed by our theoretical results.
3.1 Reward function estimation using neural network
For estimating the latent reward function, we use a fully connected neural network (NN) with depth $D\geq 2$, a hidden layer width $w$, and ReLU activations as done in Zhou et al. (2020), Zhang et al. (2021), and Verma et al. (2025). Let $h(x;\theta)$ be the output of a full-connected NN with parameters $\theta$ for context-arm feature vector $x=\varphi(c,a)$ of context $c$ and arm $a$, which we define as follows:
$h(x;\theta)=\bm{W}_{D}\text{ReLU}\left(\bm{W}_{D-1}\text{ReLU}\left(\cdots\text{ReLU}\left(\bm{W}_{1}x\right)\right)\right),$
where $\text{ReLU}(v)=\max\{v,0\}$, $\bm{W}_{1}\in\mathbb{R}^{w\times d}$, $\bm{W}_{l}\in\mathbb{R}^{w\times w}$ for $2\leq l<D$, $\bm{W}_{D}\in\mathbb{R}^{w\times 1}$. The parameters of the NN are represented by $\theta=\left(\text{vec}\left(\bm{W}_{1}\right);\cdots\text{vec}\left(\bm{W}_{D}\right)\right)$, where $\text{vec}\left(A\right)$ transforms an $m\times n$ matrix $A$ into a vector of dimension $mn$. We use $p$ to represent the total number of NN parameters, which is given by $p=dw+w^{2}(L-1)+w$, and $g(x;\theta)$ to denote the gradient of NN $h(x;\theta)$ with respect to $\theta$. At the end of each iteration $t$, the available preference dataset $\mathcal{D}_{t}=\left\{(c_{s},a_{s,w},a_{s,l},y_{s})\right\}_{s=1}^{t}$ is used to estimate the reward function $f$ by training an NN $h$ (parameterized by $\theta_{t+1}$) using gradient descent to minimize the following binary cross entropy loss function:
$\min_{\theta}\mathcal{L}_{t}(\theta)$ $=-\frac{1}{w}\sum^{t}_{s=1}\Big{[}\log\mu\big{(}h(\varphi(c_{s},a_{s,w});\theta)\ -h(\varphi(c_{s},a_{s,l});\theta)\big{)}\Big{]}+\frac{1}{2}\lambda\left\|\theta-\theta_{0}\right\|^{2}_{2},$ (2)
where $\theta_{0}$ denotes the initial parameter of the NN that is initialized according to the standard practice in neural bandits (Zhou et al., 2020; Zhang et al., 2021). Minimizing the first term in the above loss function (that involves the summation over the $t$ terms) corresponds to finding the maximum log-likelihood estimate of the parameters $\theta$.
3.2 Neural-ADB
In this section, we propose a simple yet principled and practical algorithm, Neural-ADB, that consists of two key components: Context selection and arm selection. Neural-ADB works as follows: At the beginning of the iteration $t$, we first select the context by adapting context selection strategy [^3] Note that selecting contexts uniformly at random suffer a constant suboptimality gap (Das et al., 2024, Theorem 3.2). from Das et al. (2024) as follows:
$c_{t}=\arg\!\max_{c\in\mathcal{C}}\max_{(a,b)\in\mathcal{A}\times\mathcal{A}}\left\|\varphi(c,a)-\varphi(c,b)\right\|_{V_{t}^{-1}},$ (3)
where $V_{t-1}=\frac{\lambda}{\kappa_{\mu}}\mathbb{I}_{p}+\sum_{s=1}^{t-1}z_{s}{z_{s}}^{\top}\frac{1}{w}$ in which $z_{s}=\varphi^{\prime}(c_{s},a_{s,w})-\varphi^{\prime}(c_{s},a_{s,l})=g(\varphi(c_{s},a_{s,w});\theta_{0})-g(\varphi(c_{s},a_{s,l});\theta_{0})$, and $g(\varphi(c_{s},a_{s,i});\theta_{0})/\sqrt{w}$ is used as the Random features approximation for the context-arm feature vector $\varphi(c_{s},a_{s,i})$. After selecting context $c_{t}$, Neural-ADB uses the trained NN (as an estimate for the unknown reward function) to decide which two arms must be selected. To do so, Neural-ADB uses UCB- and TS-based arm selection strategies, which efficiently balance the exploration-exploitation trade-off (Lattimore & Szepesvári, 2020) due to the bandit nature of preference feedback, as it is only observed for the selected arms.
Neural-ADB Neural Active Dueling Bandit algorithm
1:  Input parameters: $\delta\in(0,1)$, $\lambda>0$, and $w>0$
2:  Initialize: NN parameters $\theta_{1}$ and $D_{0}=\emptyset$
3:  for $t=1,\ldots,T$ do
4:     Select a context $c_{t}$ from $\mathcal{C}$ using Eq. 3
5:     Select first arm $a_{t,1}$ using Eq. 4
6:     Select second arm $a_{t,2}$ using Eq. 5 (for UCB-based arm selection) or Eq. 7 (for TS-based arm selection)
7:     Observe preference feedback $y_{t}=\mathds{1}_{\left\{a_{t,1}\succ a_{t,2}\right\}}$
8:     Update $D_{t}=D_{t-1}\cup\{(c_{t},a_{t,1},a_{t,2},y_{t})\}$
9:     Retrain NN parameters $\theta_{t+1}$ using $D_{t}$ by minimizing the loss function defined in Eq. 2
10:  end for
11:  Return policy $\pi(c)=\arg\!\max\limits_{a\in\mathcal{A}}h(\varphi(c,a);\theta_{T}),\ \forall c\in\mathcal{C}$
UCB-based arm selection strategy.   Algorithms based on Upper confidence bound (UCB) are commonly used to address the exploration-exploitation trade-off in many sequential decision-making problems (Auer et al., 2002; Abbasi-Yadkori et al., 2011; Zhou et al., 2020; Bengs et al., 2022). Our UCB-based arm selection strategy works as follows: In the iteration $t$, it selects the first arm greedily (i.e., by maximizing the output of the trained NN with parameters $\theta_{t}$) for the selected context $c_{t}$, ensuring the best-performing arm is always selected as follows:
$a_{t,1}=\arg\!\max_{a\in\mathcal{A}}h(\varphi(c_{t},a);\theta_{t}).$ (4)
The second arm $a_{t,2}$ is selected optimistically to focus on exploration, i.e., by maximizing the UCB value as follows:
$a_{t,2}=\mathop{\mathrm{argmax}}_{b\in\mathcal{A}\setminus\{a_{t,1}\}}\left[h(\varphi(c_{t},b);\theta_{t})+\text{cf}(t,c_{t},a_{t,1},b)\right],$ (5)
where $\text{cf}(t,c_{t},a_{t,1},b)=\nu_{T}\sigma_{t-1}(c_{t},a_{t,1},b)$, $\nu_{T}\doteq(\beta_{T}+B\sqrt{\lambda/\kappa_{\mu}}+1)\sqrt{\kappa_{\mu}/\lambda}$ in which $\beta_{T}\doteq\frac{1}{\kappa_{\mu}}\sqrt{\widetilde{d}+2\log(1/\delta)}$, $\widetilde{d}$ is the effective dimension (defined in Eq. 9), and
$\sigma_{t-1}^{2}(c,a,b)\doteq\frac{\lambda}{\kappa_{\mu}}\left\|\frac{1}{\sqrt{w}}(\varphi^{\prime}(c,a)-\varphi^{\prime}(c,b))\right\|^{2}_{V_{t-1}^{-1}}.$ (6)
A larger value of $\sigma_{t-1}^{2}(c_{t},a_{t,1},b)$ implies that arm $b$ is significantly different from $a_{t,1}$, given the contexts and arm pairs already selected. As a result, the second term in Eq. 5 makes the second arm different from the first arm.
TS-based arm selection strategy.   Thompson sampling (TS) selects an arm based on its probability of being the best (Agrawal & Goyal, 2013; Li et al., 2024). Several works (Li et al., 2024; Chowdhury & Gopalan, 2017; Agrawal & Goyal, 2013; Chapelle & Li, 2011) have shown that TS empirically outperforms UCB-based bandit algorithms. Therefore, we propose a TS-based arm selection strategy in which the first arm is also selected using Eq. 4 and only the second arm is selected differently. To select the second arm, it first samples a score $s_{t}(b)\sim\mathcal{N}\big{(}h(\varphi(c_{t},b);\theta_{t})-h(\varphi(c_{t},a_{t,1});\theta_{t}),\nu_{T}^{2}\sigma_{t-1}^{2}(c_{t},a_{t,1},b)\big{)}$ for every arm $b\in\mathcal{A}\setminus\{a_{t,1}\}$ and then selects the second arm as follows:
$a_{t,2}=\mathop{\mathrm{argmax}}\nolimits_{b\in\mathcal{A}\setminus\{a_{t,1}\}}s_{t}(b).$ (7)
After selecting context and arms in iteration $t$, stochastic preference feedback is observed, denoted by $y_{s}=\mathds{1}_{\left\{a_{t,1}\succ a_{t,2}\right\}}$, which is equal to $1$ if arm $a_{s,1}$ is preferred over arm $a_{s,2}$ for context $c_{t}$ and 00 otherwise. With the new observation, the preference dataset is updated to $\mathcal{D}_{t}=\mathcal{D}_{t-1}\cup\{(c_{t},a_{t,w},a_{t,l},y_{t})\}$ and then the NN is retrained using the updated preference dataset $\mathcal{D}_{t}$. Once the preference data collection process concludes (i.e., end of iteration $T$, which may not be fixed a priori), Neural-ADB returns the following policy:
$\forall c\in\mathcal{C}:\ \pi(c)=\arg\!\max_{a\in\mathcal{A}}h(\varphi(c,a);\theta_{T}).$ (8)
3.3 Theoretical Results
Let the number of arms in $\mathcal{A}$ be finite, and define $\mathbb{V}=\sum_{s=1}^{T}\sum_{(a,b)\in\mathcal{A}\times\mathcal{A}}z_{a,b}(s)z_{a,b}(s)^{\top}\frac{1}{w}$, where $z_{a,b}(s)=\varphi(c_{s},a)-\varphi(c_{s},b)$ and $C^{|\mathcal{A}|}_{2}$ denotes all pairwise combinations of arms. Then, the effective dimension of context-arm feature vectors is defined as follows:
$\widetilde{d}=\log\det\left(\frac{\kappa_{\mu}}{\lambda}\mathbb{V}+\mathbb{I}_{p}\right).$ (9)
In the following, we present a novel theoretical result that gives an upper bound on the maximum Mahalanobis norm of a vector selected from the fixed input space, measured with respect to the inverse of a positive definite Gram matrix constructed from finite, adapted samples of the same space.
Theorem 1.
Let $\{Z_{s}=z_{s}z_{s}^{\top}\}_{s=1}^{T}$ be a finite adapted sequence of self-adjoint matrices in $\mathbb{R}^{d}$. Define $\mathbb{E}\left[z_{s}z_{s}^{\top}\right]=\Sigma_{s}\leq\Sigma_{\max}$, $V_{0}=\lambda\mathbb{I}_{d}$, $V_{T}=\lambda\mathbb{I}_{d}+\sum_{s=1}^{T}z_{s}z_{s}^{\top}$. Assume $\left\|z_{s}\right\|_{2}\leq L$ for all $z\in\mathcal{Z}\subset\mathbb{R}^{d}$, $\lambda_{\min}(A)$ denote the minimum eigenvalue of a matrix $A$, and $\forall s\leq T:\ \left\|V_{s}-V_{s-1}\right\|^{2}\leq C_{s}$, where $\left\|V\right\|$ denotes the operator norm. Then, with a probability at least $1-\delta$, $\max_{z\in\mathcal{Z}}\left\|z\right\|_{V_{T}^{-1}}\leq L/G_{T},$ where $G_{T}=\sqrt{T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\nicefrac{{d}}{{\delta}}\right)}}$.
Proof sketch.   To derive the upper bound, we use various results related to the positive definite matrix (detailed in 1 of the supplementary material). First, if $V_{T}$ is a positive definite matrix $V_{T}$, then for any $z\in\mathcal{Z}$, $\left\|z\right\|_{V_{T}^{-1}}\leq\left\|z\right\|_{2}\sqrt{\lambda_{\max}(V_{n}^{-1})}=\left\|z\right\|_{2}/\sqrt{\lambda_{\min}(V_{n})}$. Thus, $\max\limits_{z\in\mathcal{Z}}\left\|z\right\|_{V_{T}^{-1}}\leq\left\|z\right\|_{2}/\sqrt{\lambda_{\min}(V_{n})}\leq L/\sqrt{\lambda_{\min}(V_{n})}$. Since $\{Z_{s}\}_{s=1}^{T}$ is a finite adapted sequence of self-adjoint matrices (i.e., $Z_{s}$ is $\mathcal{F}_{s}$-measurable for all $s$, where $\mathcal{F}_{s}$ represents all information available up to iteration $s$), we apply the Matrix Azuma inequality (Tropp, 2012) to get a high probability lower bound on $\lambda_{\min}(V_{n})$, specifically we have shown that $\lambda_{\min}(V_{T})\geq T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\nicefrac{{d}}{{\delta}}\right)}$ holds with probability at least $1-\delta$. Using this bound, we get the desired upper bound $L/G_{T}$. The full proof of 1, along with all other missing proofs, are provided in Appendix A.
This result shows that the upper bound can be expressed in terms of the number of adapted samples used to construct the matrix $V_{T}$, and it decays at a sub-linear rate as the number of samples $(T)$ increases. Notably, this result is of independent interest, as it provides valuable insights beyond the specific application of our work. Next, we give an upper bound on the worst sub-optimality gap in terms of the upper bound on the estimation error of the reward difference between any triplet consisting of a context and two arms.
Lemma 1.
Let $D_{T}=\{x_{s},a_{s,1},a_{s,2},y_{s}\}_{s=1}^{T}$ be the preference dataset collected up to the iteration $T$ and $\hat{f}_{T}$ represent the estimate of latent reward function $f$ learned from $D_{T}$. With probability at least $1-\delta$, $\forall c\in\mathcal{C},\ a,b\in\mathcal{A}:\ {\left|\left[f(\varphi(c,a))-f\left(\varphi(c,b)\right)\right]-\left[\hat{f}_{T}(\varphi(c,a))-\hat{f}_{T}(\varphi(c,b))\right]\right|}$ $\leq\beta_{T}(c,a,b)$. If $a^{\star}=\arg\!\max_{a\in\mathcal{A}}f(\varphi(c,a))$ and $\pi(c)$ is the arm selected by policy for context $c$, then, with a probability at least $1-\delta$, the worst sub-optimality gap for a policy that greedily selects an arm for a given context is upper bounded by: $\Delta_{T}^{\pi}\leq\max\limits_{c\in\mathcal{C}}\beta_{T}(c,a^{\star},\pi(c)).$
The proof follows by starting with the worst sub-optimality gap definition in Eq. 1 and then applying a series of algebraic manipulations to derive the stated result. Our next results give an upper bound on $\beta_{T}(c,a,b)$ when Neural-ADB uses different arm selection strategies.
Lemma 2.
Let $\nu_{T}=(\beta_{T}+B\sqrt{\lambda/\kappa_{\mu}}+1)\sqrt{\kappa_{\mu}/\lambda}$, where $\beta_{T}=\left(\nicefrac{{1}}{{\kappa_{\mu}}}\right)\sqrt{\widetilde{d}+2\log(1/\delta)}$ and $\delta\in(0,1)$. If $w\geq\text{poly}(T,L,K,1/\kappa_{\mu},L_{\mu},1/\lambda_{0},1/\lambda,\log(1/\delta))$, then, with a probability of at least $1-\delta$, for Neural-ADB with
  1. 1.
    UCB-based arm selection strategy, for all $c\in\mathcal{C}:$
    $\beta_{T}(c,a,b)=\nu_{T}\sigma_{T}(c,a^{\star},\pi(c))+2\varepsilon^{\prime}_{w,T},$
  2. 2.
    TS-based arm selection strategy, for all $c\in\mathcal{C}:$
    $\beta_{T}(c,a,b)=\nu_{T}\log\left(KT^{2}\right)\sigma_{T}(c,a^{\star},\pi(c))+2\varepsilon^{\prime}_{w,T},$
where $K$ denotes the maximum number of arms available in each iteration, and $\varepsilon^{\prime}_{w,T}=C_{2}w^{-1/6}\sqrt{\log w}L^{3}\left(\nicefrac{{T}}{{\lambda}}\right)^{4/3}$ for some absolute constant $C_{2}>0$, is the approximation error that decreases as the width of the NN $(w)$ increases.
The proof of the first part follows directly from Theorem 1 of (Verma et al., 2025) while the second follows from Lemma 10 with Eq. (27) of (Verma et al., 2025). Equipped with 1, 1, and 2, we will now provide an upper bound on the worse sub-optimality gap for a policy learned by Neural-ADB while using UCB- and TS-based arm selection strategy for a given context.
Theorem 2 (UCB).
Let the conditions in 1 and 2 hold. Then, with a probability with at least $1-\delta$, the worst sub-optimality gap of Neural-ADB when using UCB-based arm selection strategy is upper bounded by
$\Delta_{T}^{\pi}$ $\leq\left(\frac{\nu_{T}L}{G_{T}}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}+2\varepsilon^{\prime}_{w,T}=\tilde{O}\left(\sqrt{\frac{\tilde{d}}{T}}\right).$
Theorem 3 (TS).
Let the conditions in 1 and 2 hold. Then, with a probability with at least $1-\delta$, the worst sub-optimality gap of Neural-ADB when using TS-based arm selection strategy is upper bounded by
$\Delta_{T}^{\pi}$ $\leq\left(\frac{\nu_{T}L\log\left(KT^{2}\right)}{G_{T}}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}+2\varepsilon^{\prime}_{w,T}=\tilde{O}\left(\sqrt{\frac{\tilde{d}}{T}}\right).$
The proof follows by applying 2, setting $z^{\prime}=\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))$ in Eq. 6, and then using 1. Note that $\varepsilon^{\prime}_{w,T}=O(1/T)$ and $\widetilde{d}=\widetilde{o}(\sqrt{T})$ as long as the NN width $w$ is large enough (Verma et al., 2025). Above 2 and 3 show that the worst sub-optimality gap of the policy learned by Neural-ADB with UCB- and TS-based arm selection strategies decreases at a sub-linear rate with respect to the size of preference dataset, specifically at rate of $\tilde{O}((\tilde{d}/T)^{\frac{1}{2}})$, where $\tilde{O}$ hides the logarithmic factors and constants. Further, the decay rate of the worst sub-optimality gap for Neural-ADB improves by a factor of $\tilde{O}(({\tilde{d}\log T})^{\frac{1}{2}})$ compared to exiting algorithms (Mehta et al., 2023; Das et al., 2024), thereby bridging the gap between theory and practice.
3.4 Active Dueling Bandits with Regret Minimization
We start by defining the cumulative regret (or ‘regret’ for brevity) of a policy. After receiving preference feedback for $T$ pairs of arms, the regret of a sequential arm selection policy is given by: $\mathfrak{R}_{T}=\sum_{t=1}^{T}\left[f(\varphi(c_{t},a_{t}^{\star}))-\left({f(\varphi(c_{t},a_{t,1}))+f(\varphi(a_{t,2}))}\right)/{2}\right],$ where $a_{t}^{\star}=\arg\!\max_{a\in\mathcal{A}_{t}}f(\varphi(c_{t},a))$ denotes the arm that maximizes the reward function for a given context $c_{t}$.
In many real-world applications, such as medical treatment design (Lai & Robbins, 1985; Bengs et al., 2021) and content moderation (Avadhanula et al., 2022), both actively selecting arms and minimizing regret is required. For instance, in personalized medical treatment, active learning is used to identify the most informative treatments to test, while cumulative regret minimization ensures the system continually adapts to deliver better patient outcomes.
Since the arm selection strategies in Neural-ADB are directly adapted from UCB- and TS-based algorithms for contextual dueling bandits of Verma et al. (2025), the regret upper bounds for these algorithms also apply to Neural-ADB. For completeness, we state the regret upper bounds of Neural-ADB as follows.
Corollary 1 (UCB).
(Verma et al., 2025, Theorem 2) Let $\lambda>\kappa_{\mu}$ and $w\geq\text{poly}(T,L,K,1/\kappa_{\mu},L_{\mu},1/\lambda_{0},1/\lambda,\log(1/\delta))$. Then, with a probability of at least $1-\delta$, the regret of Neural-ADB when using UCB-based arm selection strategy is upper bounded by
$\mathfrak{R}_{T}=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}}\right).$
Corollary 2 (TS).
(Verma et al., 2025, Theorem 3) Let the conditions in 1. Then, with a probability of at least $1-\delta$, the regret of Neural-ADB when using TS-based arm selection strategy is upper bounded by
$\mathfrak{R}_{T}=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}}\right).$
Ignoring logarithmic factors and constants, the asymptotic growth rates of Neural-ADB with UCB- and TS-based arm selection strategy are identical and sub-linear.
4 Experiments
To validate our theoretical results, we empirically evaluate the performance of our algorithms on different problem instances of synthetic datasets. Specifically, we use two commonly used synthetic functions adopted from existing works on neural bandits (Zhou et al., 2020; Zhang et al., 2021; Dai et al., 2023; Verma et al., 2025): $f(x)=10(x^{\top}\theta)^{2}$ (Square), $f(x)=2\sin(x^{\top}\theta)$ (Sine), and $f(x)=\cos(3x^{\top}\theta)$ (Cosine). All experiments are repeated 10 times, and we report the average worst suboptimality gap with 95% confidence intervals (depicted as vertical lines on each curve). Additional results are also provided in Appendix A due to space constraints.
Synthetic dataset.   Let $x_{t,a}$ be the context-arm feature vector for context $c_{t}$ and an arm $a$. For all $t\geq 1$, $x_{t,a}$ is sampled uniformly at random from $(-1,1)^{d}$. We keep the number of arms constant across all rounds, denoted by $K$. In our experiments, the binary preference feedback indicating whether $x_{t,1}$ preferred over $x_{t,2}$ (representing human preference feedback) is sampled from a Bernoulli distribution with parameter $\mu\left(f(x_{t,1})-f(x_{t,2})\right)$).
Reward function estimation.   We use a neural network with $2$ hidden layers with width $50$ to estimate the latent reward function, $\lambda=1.0$, $\delta=0.05$, $d=20$, $K=10$, and fixed value of $\nu_{T}=\nu=1.0$ in all our experiments (unless we specifically indicate $d$ and $K$). Note that we did not perform any hyperparameter search for Neural-ADB.
(a) Suboptimality Gap
(b) MAE
(c) Average Regret
(d) Suboptimality Gap
(e) MAE
(f) Average Regret
(g) Suboptimality Gap
(h) MAE
(i) Average Regret
Figure 1: Performance comparison under different synthetic functions, i.e., Square function (first row), Sine function (second row), and Cosine function (third row).
Comparison with baselines.   We compare the worst suboptimality gap (defined in Eq. 1), MAE (average suboptimality gap), and average regret (defined in Section 3.4) against the different baselines of active contextual dueling bandits to evaluate the performance of UCB- and TS-variant of Neural-ADB. We use three baselines: Random, AE-Borda (Mehta et al., 2023), APO (Das et al., 2024), AE-DPO (Mehta et al., 2023) and the neural variants of AE-Borda and AE-DPO in which we use a Neural network to estimate the latent reward function. They are named AE-Borda (NN) and Neural-ADB (APO) respectively. Experimental results in Fig. 1 show that our approach – Neural-ADB (UCB) – outperforms other baselines in almost all synthetic functions (i.e., square function and sine function) in terms of the suboptimality gap. Moreover, our approach outperforms other baselines on all synthetic functions in MAE and average regret. We have included more comparisons of our approach with other baselines in other settings (e.g., different $d$ or $K$) in Appendix B.
5 Related Work
Finite-Armed Dueling Bandits.   Learning from preference feedback has been extensively studied in the bandit literature. In the finite-armed dueling bandits setting, the learner aims to find the best arm while only observing preference feedback for two selected arms (Yue & Joachims, 2009; 2011; Yue et al., 2012). To determine the best arm in dueling bandits, different criteria, such as the Borda winner, Condorcet winner, Copeland winner, or von Neumann winner, have been used while focusing on minimizing regret using only pairwise preference feedback (Zoghi et al., 2014b; Ailon et al., 2014; Zoghi et al., 2014a; Komiyama et al., 2015; Gajane et al., 2015; Saha & Gopalan, 2018; 2019a; 2019b; Verma et al., 2019; 2020a; 2020b; Zhu et al., 2023). For a comprehensive overview of algorithms for various dueling bandits settings, we refer readers to the survey by Bengs et al. (2021).
Contextual Dueling Bandits.   Many real-life applications, such as online recommendations, content moderation, medical treatment design, prompt optimization, and aligning large language models, can be effectively modeled using contextual dueling bandits, where a learner observes a context (additional information before selecting a pair of arms) and then selects the arms based on that context and observes preference feedback for the selected arms. Since the number of context-arm pairs can be potentially large or even infinite, the mean latent reward of each context-arm is assumed to be parameterized by an unknown function of its features. Common assumptions include linear reward models (Saha, 2021; Saha & Krishnamurthy, 2022; Bengs et al., 2022; Di et al., 2023; Li et al., 2024) and non-linear models (Verma et al., 2025). For our setting, we adopt the neural contextual dueling bandit algorithms proposed by Verma et al. (2025) to construct confidence ellipsoids for the latent non-linear reward function. Note that Neural-ADB can incorporate alternative confidence ellipsoids by appropriately modifying 2. Furthermore, our work addresses an active learning problem and analyzes the convergence rate of the worst sub-optimality gap, whereas Verma et al. (2025) focus on a regret minimization setting and derive upper bounds on cumulative regret.
Active contextual dueling bandits.   The work most closely related to ours is active contextual dueling bandit (Mehta et al., 2023; Das et al., 2024), which takes a principled approach to actively collecting preference datasets. However, two key differences exist between our work and existing research: the non-linear reward function and the arm-selection strategy. Existing studies typically assume a linear reward function, which may not be suitable for many real-world applications. Our work addresses this gap by extending the existing framework to incorporate non-linear reward functions in contextual dueling bandits. Additionally, existing approaches use different methods for selecting the pair of arms, leading to distinct arm-selection strategies compared to ours. As a result of these differences in both the arm-selection strategy and the non-linear reward function (which we estimate using a neural network), our analysis diverges significantly from that of prior work.
6 Conclusion
This paper studies the problem of active human preference feedback collection by modeling it as an active neural contextual dueling bandit problem. We propose Neural-ADB, a principled and practical algorithm designed for efficiently gathering human preference feedback in scenarios where the reward function is non-linear. Exploiting the neural contextual dueling bandit framework, Neural-ADB extends its applicability to a broad range of real-world applications, including online recommendation systems and LLM alignment. Our theoretical analysis demonstrates that the worst suboptimality gap of Neural-ADB decays at a sub-linear rate as the preference dataset grows. Finally, our experimental results further validate these theoretical findings. An interesting direction for future work is applying Neural-ADB to real-life applications such as LLM alignment (Mehta et al., 2023; Das et al., 2024). From a theoretical perspective, exploring the non-stationary setting presents another promising future direction.
Acknowledgements
This research is supported by the National Research Foundation (NRF), Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. The Mens, Manus, and Machina (M3S) is an interdisciplinary research group (IRG) of the Singapore MIT Alliance for Research and Technology (SMART) centre.
Impact Statement
This paper’s primary contributions are theoretical, so we do not anticipate any immediate negative societal impacts. However, our algorithms could be used to improve online DPO/RLHF for efficient LLM alignment, which would benefit society. While the potential negative societal impacts associated with LLM alignment may also extend to our work, mitigation strategies for preventing the misuse of LLM alignment would also help safeguard against any potential misuse of our algorithms.
Appendix A Leftover Proofs
To simplify the presentation, we use a common error probability of $\delta$ for all probabilistic statements. Our final results naturally follow by applying a union bound over all individual $\delta$. Next, we will describe the key properties of positive definite matrices crucial for the subsequent proofs. These properties form the basis for several key parts of our analysis.
Fact 1 (Properties of a positive definite matrix).
Let $V_{0}=\lambda\mathbb{I}_{d}$, $V_{T}=V_{0}+\sum_{s=1}^{T}z_{s}z_{s}^{\top}$ be a positive definite matrix, where $\lambda>0$, $z_{s}\in\mathbb{R}^{d}$, and $\{Z_{s}=z_{s}z_{s}^{\top}\}_{s=1}^{T}$ is a finite adapted sequence of self-adjoint matrices, i.e., $V_{s}$ and $Z_{s}$ are $\mathcal{F}_{s}$-measurable for all $s$, where $\mathcal{F}_{s}$ represents all information available up to $s$. We use $\lambda_{\max}(V_{T})$ and $\lambda_{\min}(V_{T})$ to denote the maximum and minimum eigenvalue of matrix $V_{T}$. Then, the following properties hold for $V_{T}$:
  1. 1.
    $\lambda_{\max}(V_{T})\leq\lambda+\sum_{s=1}^{T}\lambda_{\max}\left(z_{s}z_{s}^{\top}\right)$.
  2. 2.
    $\lambda_{\min}(V_{T})\geq\lambda+\sum_{s=1}^{T}\lambda_{\min}\left(z_{s}z_{s}^{\top}\right)$.
  3. 3.
    $\lambda_{\max}\left(z_{s}z_{s}^{\top}\right)\leq\left\|z_{s}\right\|_{2}^{2}\ \implies\ \lambda_{\max}(V_{T})\leq\lambda+\sum_{s=1}^{T}\left\|z_{s}\right\|_{2}^{2}$.
  4. 4.
    $\lambda_{\max}(V_{T}^{-1})={1}/{\lambda_{\min}(V_{T})}$.
  5. 5.
    $\lambda_{\max}(V_{T})=-{\lambda_{\min}(-V_{T})}$.
  6. 6.
    Let $\lambda_{i}(V)$ be the $i$-th eigenvalue of matrix $V$. If $W$ is any Hermitian matrix, then, from Weyl’s inequality:
    $1.\ \lambda_{i}(V_{T})+\lambda_{\min}(W)\leq\lambda_{i}(V_{T}+W)\leq\lambda_{i}(V_{T})+\lambda_{\max}(W)\text{ and }$
    $2.\ \lambda_{i}(V_{T})-\lambda_{\max}(W)\leq\lambda_{i}(V_{T}-W)\leq\lambda_{i}(V_{T})-\lambda_{\min}(W).$
  7. 7.
    For any vector $z\in\mathbb{R}^{d}$, the Mahalanobis norm of $z$ with respect to $V_{T}$ is $\left\|z\right\|_{V_{T}}=\sqrt{z^{\top}V_{T}z}\leq\left\|z\right\|_{2}\sqrt{\lambda_{\max}(V_{T})}$.
  8. 8.
    For $a>0:\ \left\|az\right\|_{V_{T}}=a\left\|z\right\|_{V_{T}}$ and $\lambda_{i}(aV_{T})=a\lambda_{i}(V_{T})$.
  9. 9.
    Let $z_{\max}=\arg\!\max_{z\in\mathbb{R}^{d}}\left\|z\right\|_{V_{T}}$. Then, $z_{\max}$ is the eigenvector of $V_{T}$ corresponding to its largest eigenvalue.
  10. 10.
    Let $\forall z\in\mathbb{R}^{d}\colon\ \left\|z\right\|_{2}\leq L$. Then, $\max_{z\in\mathbb{R}^{d}}\left\|z\right\|_{V_{T}^{-1}}\leq\left\|z\right\|_{2}\sqrt{\lambda_{\max}(V_{T}^{-1})}\leq L/\sqrt{\lambda_{\min}(V_{T})}$.
  11. 11.
    Let $\delta\in(0,1)$, $\forall s\leq T:\ \left\|V_{s}-V_{s-1}\right\|^{2}\leq C_{s}$, where $\left\|A\right\|$ denotes the operator norm. Then, using Theorem 7.1 and Corollary 7.2 of Tropp (2012), with probability at least $1-\delta$,
    $\mathbb{P}\left\{\lambda_{\max}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)\geq\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\frac{d}{\delta}\right)}\right\}\leq\delta.$
A.1 Proof of 1
We now prove the upper bound on the maximum Mahalanobis norm of a vector from the fixed input space, measured with respect to the inverse of a positive definite Gram matrix defined by finite, adapted samples from the same input space.
See 1
Proof.
Using Property 11 from 1 with $Y_{T}-\mathbb{E}\left[Y_{T}\right]=\mathbb{E}\left[V_{T}\right]-V_{T}$, we have
$\mathbb{P}\left\{\lambda_{\max}\left(\mathbb{E}\left[V_{T}\right]-V_{T}\right)\geq\tau\right\}\leq d\exp\left(\frac{-\tau^{2}}{8\sum_{s=1}^{T}C_{s}}\right)$
[…]
$\implies$ $\mathbb{P}\left\{\lambda_{\min}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)\leq-\tau\right\}\leq d\exp\left(\frac{-\tau^{2}}{8\sum_{s=1}^{T}C_{s}}\right).$
Using upper bound on $\lambda_{\min}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)$ from Property 6 in 1, we get
$\implies$ $\mathbb{P}\left\{\lambda_{\min}(V_{T})-\lambda_{\min}(\mathbb{E}\left[V_{T}\right])\leq-\tau\right\}\leq d\exp\left(\frac{-\tau^{2}}{8\sum_{s=1}^{T}C_{s}}\right)$
$\implies$ $\mathbb{P}\left\{\lambda_{\min}(V_{T})\leq\lambda_{\min}(\mathbb{E}\left[V_{T}\right])-\tau\right\}\leq d\exp\left(\frac{-\tau^{2}}{8\sum_{s=1}^{T}C_{s}}\right).$
Note that $\mathbb{E}\left[V_{T}\right]=\mathbb{E}\left[\sum_{t=1}^{T}z_{s}z_{s}^{\top}\right]=\sum_{t=1}^{T}\mathbb{E}\left[z_{s}z_{s}^{\top}\right]=\sum_{t=1}^{T}\Sigma_{s}\leq T\Sigma_{\max}$. Thus, we get
$\mathbb{P}\left\{\lambda_{\min}(V_{T})\leq T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\frac{d}{\delta}\right)}\right\}\leq\delta.$
Therefore, with probability at least $1-\delta$, $\lambda_{\min}(V_{T})\geq T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\nicefrac{{d}}{{\delta}}\right)}$. Using Property 10 from 1, we now use to prove our key result as follows:
$\max_{z\in\mathcal{Z}}\left\|z\right\|_{V_{T}^{-1}}$ $\leq L/\sqrt{\lambda_{\min}(V_{T})}$
$\leq L/\sqrt{T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\frac{d}{\delta}\right)}}$
$=L/G_{T}$
$\implies\max_{z\in\mathcal{Z}}\left\|z\right\|_{V_{T}^{-1}}$ $\leq=L/G_{T}.\qed$
A.2 Proof of 1
Our next results gives an upper bound of worst sub-optimality gap in terms of the upper bound of estimation error in the reward difference between any triple of context and two arms. See 1
Proof.
Define $a^{\star}=\arg\!\max\limits_{a\in\mathcal{A}}f(\varphi(c,a))$. Recall the definition of worst suboptimality across all contexts, which is :
$\Delta_{\mathcal{D}_{T}}^{\pi}$ $=\max_{c\in\mathcal{C}}\left[\max_{a\in\mathcal{A}}f(\varphi(c,a))-f\left(\varphi(c,\pi(c))\right)\right]$
$=\max_{c\in\mathcal{C}}\left[f(\varphi(c,a^{\star}))-f\left(\varphi(c,\pi(c))\right)\right]$
$=\max_{c\in\mathcal{C}}\left[f(\varphi(c,a^{\star}))-f\left(\varphi(c,\pi(c))\right)+\hat{f}_{T}(\varphi(c,a^{\star}))-\hat{f}_{T}(\varphi(c,a^{\star}))\right]$
$\leq\max_{c\in\mathcal{C}}\left|\left[f(\varphi(c,a^{\star}))-f\left(\varphi(c,\pi(c))\right)\right]+\left[\hat{f}_{T}(\varphi(c,\pi(c)))-\hat{f}_{T}(\varphi(c,a^{\star}))\right]\right|$
$=\max_{c\in\mathcal{C}}\left|\left[f(\varphi(c,a^{\star}))-f\left(\varphi(c,\pi(c))\right)\right]-\left[\hat{f}_{T}(\varphi(c,a^{\star}))-\hat{f}_{T}(\varphi(c,\pi(c)))\right]\right|$
$\implies\Delta_{\mathcal{D}_{T}}^{\pi}$ $\leq\max_{c\in\mathcal{C}}\beta_{T}(c,a^{\star},\pi(c)).$
The inequality follows from the fact we have greedy policy, i.e., $\pi(c)=\arg\!\min_{a\in\mathcal{A}}\hat{f}_{T}(\varphi(c,a))$ for any context $c$. Therefore, if $\pi(c)\neq a^{\star}$, then $\hat{f}_{T}(\varphi(c,\pi(c)))\geq\hat{f}_{T}(\varphi(c,a^{\star}))$ must hold. ∎
A.3 Proof of 2 and 3
Equipped with 1, 1, and 2, we will next prove the upper bound on the worst sub-optimality gap for a policy learned by Neural-ADB while using UCB- and TS-based arm selection strategy for a given context.
See 2
Proof.
Using 2 and setting value of $\beta_{T}(c,a^{\star},\pi(c))$ using 2 and Eq. 6, we have
[…]
As $\nu_{T}$ and $\varepsilon^{\prime}_{w,T}$ independent of context $c$, we get
$\Delta_{T}^{\pi}$ $\leq\nu_{T}\max_{c\in\mathcal{C}}\left(\sigma_{T}(c,a^{\star},\pi(c))\right)+2\varepsilon^{\prime}_{w,T}$
[…]
$\leq\tilde{O}\left(\sqrt{\frac{\tilde{d}}{T}}\right).$
See 3
Proof.
Using 2 and setting value of $\beta_{T}(c,a^{\star},\pi(c))$ using 2 and Eq. 6, we have
[…]
The value of $\nu_{T}$ and $\varepsilon^{\prime}_{w,T}$ independent of context $c$. By following similar steps to those in the proof of 2, we have
$\Delta_{T}^{\pi}$ $\leq\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sigma_{T}(c,a^{\star},\pi(c))\right)+2\varepsilon^{\prime}_{w,T}$
$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}}}\left\|\frac{\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))}{\sqrt{w}}\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\left\|\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
$=\nu_{T}\log\left(KT^{2}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\max_{c\in\mathcal{C}}\left(\left\|\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
$\leq\nu_{T}\log\left(KT^{2}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\left(\frac{L}{\sqrt{T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\frac{d}{\delta}\right)}}}\right)+2\varepsilon^{\prime}_{w,T}$
$\leq\tilde{O}\left(\sqrt{\frac{\tilde{d}}{T}}\right).$
Appendix B Experimental Details and Additional Results
B.1 Experimental Details
Computational resources used for experiments.   All experiments were conducted on a server equipped with an AMD EPYC 7543 32-Core Processor, 256GB of RAM, and 8 NVIDIA GeForce RTX 3080 GPUs.
Practical considerations.   Based on the neural tangent kernel (NTK) theory (Jacot et al., 2018), the initial gradient $g(x;\theta_{0})$ can be used as the original feature vector $x$ as $g(x;\theta_{0})$ effectively represents the random Fourier features of the NTK. To make our algorithm more practical, we use common practices in neural bandits (Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2025). Specifically, we replaced the theoretical regularization parameter $\frac{1}{2}w\lambda\left\|\theta-\theta_{0}\right\|^{2}_{2}$ (where $w$ is the NN’s width) with the simpler $\lambda\left\|\theta\right\|^{2}_{2}$ in the loss function (defined in Eq. 2) that is used to train our NN. We retrain the neural network after every $20$ rounds for $50$ gradient steps across all experiments.
B.2 Additional Results
Next, we present the additional experiment results comparing the performance of our proposed algorithms across varying input dimensions and different numbers of arms. We then evaluate the performance of Neural-ADB (TS) under various input dimensions and numbers of arms.
(a) Sub-Optimality Gap
(b) MAE
(c) Average Regret
(d) Sub-Optimality Gap
(e) MAE
(f) Average Regret
Figure 2: Performance comparison under different input dimensionalities, i.e., 20-dimensional input (first row), 40-dimensional input (second row). The arm size is 10. We are using Square function for these results.
(a) Sub-Optimality Gap
(b) MAE
(c) Average Regret
(d) Sub-Optimality Gap
(e) MAE
(f) Average Regret
(g) Sub-Optimality Gap
(h) MAE
(i) Average Regret
Figure 3: Performance comparison under different numbers of arms. The first row is when the arm size is 5, and the second row is when the arm size is 10, the third row is when the arm size is 15. The dimensionality of input is 20. We are using Square function for these results.
[…]
(a) Suboptimality Gap
(b) MAE
(c) Average Regret
(d) Suboptimality Gap
(e) MAE
(f) Average Regret
Figure 4: Performance of Neural-ADB (UCB) under different input dimensions (first row) and numbers of arms (second row). We are using Square function for these results.
(a) Sub-Optimality Gap
(b) MAE
(c) Average Regret
(d) Sub-Optimality Gap
(e) MAE
(f) Average Regret
Figure 5: Performance of Neural-ADB (TS) under different input dimensions (first row) and numbers of arms (second row). We are using Square function for these results.

### v2

@@
-Active Human Feedback Collection via Neural Contextual Dueling Bandits
+hyperrefIgnoring empty anchor
 Abstract
@@
 We theoretically show that when preference feedback follows the Bradley-Terry-Luce model, the worst sub-optimality gap of the policy learned by Neural-ADB decreases at a sub-linear rate as the preference dataset increases.
-Our experimental results on problem instances derived from synthetic preference datasets further validate the effectiveness of Neural-ADB.
+Our experimental results on preference datasets further corroborate the effectiveness of Neural-ADB.
 1 Introduction
-Collecting human preference feedback is essential in many real-life applications, like online recommendations (Kohli et al., 2013; Wu et al., 2023; Zhang & Wang, 2023; Yang et al., 2024), content moderation (Avadhanula et al., 2022), medical treatment design (Lai & Robbins, 1985; Bengs et al., 2021), prompt optimization (Lin et al., 2024), and aligning large language models (Bai et al., 2022; Mehta et al., 2023; Menick et al., 2022; Chaudhari et al., 2024; Das et al., 2024; Ji et al., 2024), to ensure systems effectively align with user preferences and exhibit desired behaviors.
+Collecting human preference feedback is essential in many real-life applications, like online recommendations (Kohli et al., 2013; Wu et al., 2023; Zhang and Wang, 2023; Yang et al., 2024), content moderation (Avadhanula et al., 2022), medical treatment design (Lai and Robbins, 1985; Bengs et al., 2021), prompt optimization (Lin et al., 2024), and aligning large language models (Bai et al., 2022; Menick et al., 2022; Mehta et al., 2023; Chaudhari et al., 2024; Das et al., 2024; Ji et al., 2024), to ensure systems effectively align with user preferences and exhibit desired behaviors.
 However, this process is often costly due to the need for skilled evaluators, the complexity of tasks, and the time-intensive nature of producing high-quality, reliable human feedback.
 To address the challenge of balancing cost and effectiveness in aligning systems, this paper proposes principled and practical algorithms for efficiently collecting human feedback sequentially and adaptively to achieve the desired system behavior.
-Specifically, this paper aims to answer the following fundamental question:
+Specifically, we aim to answer the following fundamental question:
 How to achieve desired system behavior while using as minimum human feedback as possible?
@@
 The preference feedback between two context-arm pairs is commonly assumed to follow the Bradley-Terry-Luce (BTL) model [^1] For more than two context-arm pairs, preferences are typically modeled using the Plackett-Luce model (Soufiani et al., 2014).
-(Hunter, 2004; Luce, 2005; Saha, 2021; Bengs et al., 2022; Li et al., 2024; Lin et al., 2024; Verma et al., 2025) in which the probability of preferring a context-arm pair over others is proportional to the exponential of its reward.
-In many real-life applications, the number of context-arm pairs (e.g., prompt-response pair in prompt optimization or user-movie pair in online movie recommendation) can be large or even infinite.
-Therefore, the reward for each context-arm pair is assumed to be parameterized by an unknown function of its feature vector.
+(Hunter, 2004; Bengs et al., 2022; Li et al., 2024; Lin et al., 2024; Verma et al., 2025) in which the probability of preferring a context-arm pair over others is proportional to the exponential of its reward.
+In many real-life applications, the number of context-arm pairs (e.g., user-movie pair in online movie recommendation) can be large or even infinite.
+Therefore, the reward for each context-arm pair is assumed to be an unknown function of its feature vector, such as a linear function (Mehta et al., 2023; Das et al., 2024).
 To better align the system for optimal performance, we consider two key components: context selection and arm selection.
@@
 An efficient arm selection strategy requires estimating the reward function to guide the arm selection process effectively.
-Since the reward function may not always be linear in practice, this paper parameterizes the reward function via a non-linear function, which needs to be estimated using the available preference dataset.
-This non-linear function can be estimated by using methods like Gaussian processes (Williams & Rasmussen, 2006; Srinivas et al., 2010) or neural networks (Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2025).
+Since the reward function may not always be linear in practice, this paper parameterizes the reward function via a non-linear function, which needs to be estimated from the available preference dataset by using methods like Gaussian processes (Williams and Rasmussen, 2006; Srinivas et al., 2010) or neural networks (Zhou et al., 2020; Zhang et al., 2021).
 However, Gaussian processes have limited expressive power and fail to optimize highly complex functions.
-In contrast, neural networks (NNs) have greater expressive power, making them well-suited for modeling complex functions (Dai et al., 2023; Lin et al., 2023; 2024; Verma et al., 2025).
+In contrast, neural networks (NNs) have greater expressive power, making them well-suited for modeling complex functions (Dai et al., 2023; Lin et al., 2023, 2024; Verma et al., 2025).
 In this paper, we propose a neural active contextual dueling bandit algorithm, Neural-ADB, which uses an NN to estimate the unknown reward function using the available preference dataset.
@@
 1) of the policy learned by Neural-ADB decreases at a sub-linear rate as the preference dataset size increases.
-Specifically, our contribution are as follows:
+Specifically, our key contributions can be summarized as follows:
   • We introduce the setting of active contextual dueling bandits with a non-linear reward function in Section 2.
-In Section 3, we propose a practical neural active contextual dueling bandit algorithm, Neural-ADB, which uses an NN to estimate the unknown reward function from the available preference dataset and then uses this estimate into the arm selection strategies.
-  • One of the main theoretical contributions of this paper is proving an upper bound on the maximum Mahalanobis norm of a vector from the fixed input space, as measured with respect to the inverse of a positive definite Gram matrix (1), where the gram matrix is constructed using finite, adapted samples from that input space.
+In Section 3, we propose a neural active contextual dueling bandit algorithm, Neural-ADB, which uses an NN to estimate the unknown reward function from the available preference dataset and then uses this estimate into the arm selection strategies.
+  • We prove an upper bound on the maximum Mahalanobis norm of a vector from the fixed input space, as measured with respect to the inverse of a positive definite Gram matrix (1), where the gram matrix is constructed using finite, adapted samples from that input space.
 We show that this upper bound decays at a sub-linear rate as the number of samples used in the Gram matrix increases.
-This result itself is of independent interest, as it gives valuable insights beyond the specific application of our work.
+This theoretical result itself is of independent interest, as it gives valuable insights beyond the specific application of our work.
   • We prove that the worst sub-optimality gap of the policy learned by Neural-ADB with both of our arm selection strategies (2 and 3) decreases at a sub-linear rate with respect to the size of preference dataset, specifically at rate of $\tilde{O}((\tilde{d}/T)^{\frac{1}{2}})$, where $\tilde{O}$ hides the logarithmic factors and constants, and $\tilde{d}$ is the effective dimension of context-arm feature vectors.
 The decay rate of the worst sub-optimality gap for Neural-ADB improves by a factor of $\tilde{O}(({\tilde{d}\log T})^{\frac{1}{2}})$ compared to exiting algorithms (Mehta et al., 2023; Das et al., 2024), thus bridging the gap between theory and practice.
-  • Finally, in Section 4, our experimental results further validate the different performance aspects of Neural-ADB.
+  • Finally, in Section 4, our experimental results further validate the different performance aspects of Neural-ADB, highlighting its sample efficiency for preference data collection.
 2 Problem Setting
-We formulate active human preference feedback collection as an active contextual dueling bandit problem, where a labeler (human or simulator) provides preference feedback for a chosen pair of options (arms), such as prompt-response pairs in prompt optimization (Lin et al., 2024) and LLM alignment (Mehta et al., 2023; Das et al., 2024).
-Active contextual dueling bandit problem.
+We model active human preference feedback collection as an active contextual dueling bandit problem, where a labeler (human or simulator) provides preference feedback for a chosen pair of arms.
+Active contextual dueling bandit.
 We consider an active contextual dueling bandit problem, where the underlying latent reward function can be non-linear.
@@
 After selecting the triplet of context and two arms, the learner receives a stochastic preference feedback $y_{t}$, where $y_{t}=1$ implies the arm $a_{t,1}$ is preferred over arm $a_{t,2}$ for the context $c_{t}$ and $y_{t}=0$ otherwise.
-We use $\varphi(c_{t},a)$ to denote the context-arm feature vector for context $c_{t}$ and an arm $a$, where $\varphi:\mathcal{C}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ is a known feature map.
+We use $\varphi(c_{t},a)$ to denote the context-arm feature vector for context $c_{t}$ and an arm $a$, where $\varphi:\mathcal{C}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ is a known feature map, such as one that concatenates the context and arm features.
 Preference model.
@@
 Assumption 1.
-  • Let $\kappa_{\mu}\doteq\inf\limits_{c\in\mathcal{C},a,b\in\mathcal{A}}\dot{\mu}(f(\varphi(c,a))-f(\varphi(c,b)))>0$ for all triplets of context and arms.
+  • Let $\kappa_{\mu}\doteq\inf\limits_{c\in\mathcal{C},a,b\in\mathcal{A}}\dot{\mu}(f(\varphi(c,a))-f(\varphi(c,b)))>0$ for all triplets of context $(c)$ and pair of arms $(a,b)$.
   • The link function $\mu:\mathbb{R}\rightarrow[0,1]$ is continuously differentiable and Lipschitz with constant $L_{\mu}$. For logistic function, we have $L_{\mu}\leq 1/4$.
@@
 We denote the collected preference dataset up to $T$ iterations by $\mathcal{D}_{T}=\{(c_{s},a_{s,w},a_{s,l},y_{s})\}_{s=1}^{T}$, where $a_{s,w}\succ a_{s,l}$ for the selected context $c_{s}$ in iteration $s$.
-We aim to learn a policy, $\pi:\mathcal{C}\rightarrow\mathcal{A}$ from the collected preference dataset $\mathcal{D}_{T}$ that achieves the worst suboptimality across all contexts in $\mathcal{C}$, which is defined as follows:
+We aim to learn a policy, $\pi:\mathcal{C}\rightarrow\mathcal{A}$ from the collected preference dataset $\mathcal{D}_{T}$ that achieves the worst sub-optimality gap across all contexts in $\mathcal{C}$, which is defined as follows:
 $\Delta_{\mathcal{D}_{T}}^{\pi}=\max_{c\in\mathcal{C}}\left[\max_{a\in\mathcal{A}}f(\varphi(c,a))-f\left(\varphi(c,\pi(c))\right)\right],$ (1)
@@
 3 Algorithm for Active Human Preference Feedback Collection
-This section introduces Neural-ADB, a simple yet principled and practical algorithm designed to efficiently select context-arm pairs for collecting preference feedback.
+In this section, we introduce Neural-ADB, a simple yet principled and practical algorithm designed to efficiently select context-arm pairs for collecting preference feedback.
 Neural-ADB consists of two main components: context selection and arm selection.
 Since the arm selection strategy depends on the estimated reward function, we first explain how an NN can be used to estimate the unknown reward function.
-We will then give details of the context and arm selection strategies, followed by our theoretical results.
+We will then give details of the context and arm selection strategies, followed by our theoretical results that validate the effectiveness of Neural-ADB.
 3.1 Reward function estimation using neural network
@@
 The parameters of the NN are represented by $\theta=\left(\text{vec}\left(\bm{W}_{1}\right);\cdots\text{vec}\left(\bm{W}_{D}\right)\right)$, where $\text{vec}\left(A\right)$ transforms an $m\times n$ matrix $A$ into a vector of dimension $mn$.
-We use $p$ to represent the total number of NN parameters, which is given by $p=dw+w^{2}(L-1)+w$, and $g(x;\theta)$ to denote the gradient of NN $h(x;\theta)$ with respect to $\theta$.
-At the end of each iteration $t$, the available preference dataset $\mathcal{D}_{t}=\left\{(c_{s},a_{s,w},a_{s,l},y_{s})\right\}_{s=1}^{t}$ is used to estimate the reward function $f$ by training an NN $h$ (parameterized by $\theta_{t+1}$) using gradient descent to minimize the following binary cross entropy loss function:
+We use $p$ to represent the total number of NN parameters, which is given by $p=dw+w^{2}(D-1)+w$, and $g(x;\theta)$ to denote the gradient of NN $h(x;\theta)$ with respect to $\theta$.
+At the end of each iteration $t$, the preference dataset $\mathcal{D}_{t}=\left\{(c_{s},a_{s,w},a_{s,l},y_{s})\right\}_{s=1}^{t}$ is used to estimate the reward function $f$ by training an NN $h$ (parameterized by $\theta_{t+1}$) using gradient descent to minimize the following binary cross entropy loss function:
 $\min_{\theta}\mathcal{L}_{t}(\theta)$ $=-\frac{1}{w}\sum^{t}_{s=1}\Big{[}\log\mu\big{(}h(\varphi(c_{s},a_{s,w});\theta)\ -h(\varphi(c_{s},a_{s,l});\theta)\big{)}\Big{]}+\frac{1}{2}\lambda\left\|\theta-\theta_{0}\right\|^{2}_{2},$ (2)
-where $\theta_{0}$ denotes the initial parameter of the NN that is initialized according to the standard practice in neural bandits (Zhou et al., 2020; Zhang et al., 2021).
+where $\theta_{0}$ denotes the initial parameter of the NN that is initialized according to the standard practice in neural bandits (Zhou et al., 2020; Zhang et al., 2021) (see Algorithm 1 in Zhang et al.
+(2021) for details).
 Minimizing the first term in the above loss function (that involves the summation over the $t$ terms) corresponds to finding the maximum log-likelihood estimate of the parameters $\theta$.
 3.2 Neural-ADB
-In this section, we propose a simple yet principled and practical algorithm, Neural-ADB, that consists of two key components: Context selection and arm selection.
-Neural-ADB works as follows: At the beginning of the iteration $t$, we first select the context by adapting context selection strategy [^3] Note that selecting contexts uniformly at random suffer a constant suboptimality gap (Das et al., 2024, Theorem 3.2).
+We next propose a simple yet principled and practical algorithm, Neural-ADB, that consists of two key components: Context selection and arm selection.
+Neural-ADB works as follows: At the beginning of the iteration $t$, we first select the context as follows:
+$c_{t}=\arg\!\max_{c\in\mathcal{C}}\max_{(a,b)\in\mathcal{A}\times\mathcal{A}}\left\|\varphi(c,a)-\varphi(c,b)\right\|_{V_{t}^{-1}},$ (3)
+where $V_{t-1}=\frac{\lambda}{\kappa_{\mu}}\mathbb{I}_{p}+\sum_{s=1}^{t-1}z_{s}{z_{s}}^{\top}\frac{1}{w}$ in which $z_{s}=\varphi(c_{s},a_{s,w})-\varphi(c_{s},a_{s,l})=g(\varphi(c_{s},a_{s,w});\theta_{0})-g(\varphi(c_{s},a_{s,l});\theta_{0})$, and $g(\varphi(c_{s},a_{s,i});\theta_{0})/\sqrt{w}$ is used as the Random features approximation for the context-arm feature vector $\varphi(c_{s},a_{s,i})$.
+This strategy is adapted from the context selection strategy [^3] Note that selecting contexts uniformly at random suffer a constant sub-optimality gap (Das et al., 2024, Theorem 3.2).
 from Das et al.
-(2024) as follows:
-$c_{t}=\arg\!\max_{c\in\mathcal{C}}\max_{(a,b)\in\mathcal{A}\times\mathcal{A}}\left\|\varphi(c,a)-\varphi(c,b)\right\|_{V_{t}^{-1}},$ (3)
-where $V_{t-1}=\frac{\lambda}{\kappa_{\mu}}\mathbb{I}_{p}+\sum_{s=1}^{t-1}z_{s}{z_{s}}^{\top}\frac{1}{w}$ in which $z_{s}=\varphi^{\prime}(c_{s},a_{s,w})-\varphi^{\prime}(c_{s},a_{s,l})=g(\varphi(c_{s},a_{s,w});\theta_{0})-g(\varphi(c_{s},a_{s,l});\theta_{0})$, and $g(\varphi(c_{s},a_{s,i});\theta_{0})/\sqrt{w}$ is used as the Random features approximation for the context-arm feature vector $\varphi(c_{s},a_{s,i})$.
-After selecting context $c_{t}$, Neural-ADB uses the trained NN (as an estimate for the unknown reward function) to decide which two arms must be selected.
-To do so, Neural-ADB uses UCB- and TS-based arm selection strategies, which efficiently balance the exploration-exploitation trade-off (Lattimore & Szepesvári, 2020) due to the bandit nature of preference feedback, as it is only observed for the selected arms.
+(2024).
+After selecting context $c_{t}$, Neural-ADB uses the trained NN (as an estimate of the unknown reward function) to decide which two arms must be selected.
+To do so, Neural-ADB uses UCB- and TS-based arm selection strategies, which efficiently balance the trade-off between exploration and exploitation (Lattimore and Szepesvári, 2020) due to the bandit nature of preference feedback, as preference feedback is only observed for the selected pair of arms.
 Neural-ADB Neural Active Dueling Bandit algorithm
@@
 $a_{t,1}=\arg\!\max_{a\in\mathcal{A}}h(\varphi(c_{t},a);\theta_{t}).$ (4)
-The second arm $a_{t,2}$ is selected optimistically to focus on exploration, i.e., by maximizing the UCB value as follows:
+The second arm $a_{t,2}$ is selected optimistically by maximizing the UCB value as follows:
 $a_{t,2}=\mathop{\mathrm{argmax}}_{b\in\mathcal{A}\setminus\{a_{t,1}\}}\left[h(\varphi(c_{t},b);\theta_{t})+\text{cf}(t,c_{t},a_{t,1},b)\right],$ (5)
@@
 9), and
-$\sigma_{t-1}^{2}(c,a,b)\doteq\frac{\lambda}{\kappa_{\mu}}\left\|\frac{1}{\sqrt{w}}(\varphi^{\prime}(c,a)-\varphi^{\prime}(c,b))\right\|^{2}_{V_{t-1}^{-1}}.$ (6)
+$\sigma_{t-1}^{2}(c,a,b)\doteq\frac{\lambda}{\kappa_{\mu}}\left\|\frac{1}{\sqrt{w}}(\varphi(c,a)-\varphi(c,b))\right\|^{2}_{V_{t-1}^{-1}}.$ (6)
 A larger value of $\sigma_{t-1}^{2}(c_{t},a_{t,1},b)$ implies that arm $b$ is significantly different from $a_{t,1}$, given the contexts and arm pairs already selected.
 As a result, the second term in Eq.
-5 makes the second arm different from the first arm.
+5 makes the second arm different from the first arm which ensures exploration.
 TS-based arm selection strategy.
-Thompson sampling (TS) selects an arm based on its probability of being the best (Agrawal & Goyal, 2013; Li et al., 2024).
-Several works (Li et al., 2024; Chowdhury & Gopalan, 2017; Agrawal & Goyal, 2013; Chapelle & Li, 2011) have shown that TS empirically outperforms UCB-based bandit algorithms.
-Therefore, we propose a TS-based arm selection strategy in which the first arm is also selected using Eq.
-4 and only the second arm is selected differently.
-To select the second arm, it first samples a score $s_{t}(b)\sim\mathcal{N}\big{(}h(\varphi(c_{t},b);\theta_{t})-h(\varphi(c_{t},a_{t,1});\theta_{t}),\nu_{T}^{2}\sigma_{t-1}^{2}(c_{t},a_{t,1},b)\big{)}$ for every arm $b\in\mathcal{A}\setminus\{a_{t,1}\}$ and then selects the second arm as follows:
+Thompson sampling (TS) selects an arm based on its probability of being the best (Thompson, 1933).
+Several works (Chapelle and Li, 2011; Agrawal and Goyal, 2013; Chowdhury and Gopalan, 2017; Li et al., 2024) have shown that TS empirically outperforms UCB-based bandit algorithms.
+Therefore, we also propose a TS-based arm selection strategy in which the first arm is also selected using Eq.
+4 and the second arm is selected differently.
+To select the second arm, it first samples a score $s_{t}(b)\sim\mathcal{N}\big{(}h(\varphi(c_{t},b);\theta_{t})-h(\varphi(c_{t},a_{t,1});\theta_{t}),\nu_{T}^{2}\sigma_{t-1}^{2}(c_{t},a_{t,1},b)\big{)}$ for every arm $b\in\mathcal{A}\setminus\{a_{t,1}\}$ and then selects the second arm that maximizes the samples scores as follows:
 $a_{t,2}=\mathop{\mathrm{argmax}}\nolimits_{b\in\mathcal{A}\setminus\{a_{t,1}\}}s_{t}(b).$ (7)
@@
 where $K$ denotes the maximum number of arms available in each iteration, and $\varepsilon^{\prime}_{w,T}=C_{2}w^{-1/6}\sqrt{\log w}L^{3}\left(\nicefrac{{T}}{{\lambda}}\right)^{4/3}$ for some absolute constant $C_{2}>0$, is the approximation error that decreases as the width of the NN $(w)$ increases.
-The proof of the first part follows directly from Theorem 1 of (Verma et al., 2025) while the second follows from Lemma 10 with Eq.
-(27) of (Verma et al., 2025).
 Equipped with 1, 1, and 2, we will now provide an upper bound on the worse sub-optimality gap for a policy learned by Neural-ADB while using UCB- and TS-based arm selection strategy for a given context.
@@
 $\Delta_{T}^{\pi}$ $\leq\left(\frac{\nu_{T}L\log\left(KT^{2}\right)}{G_{T}}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}+2\varepsilon^{\prime}_{w,T}=\tilde{O}\left(\sqrt{\frac{\tilde{d}}{T}}\right).$
-The proof follows by applying 2, setting $z^{\prime}=\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))$ in Eq.
+The proof follows by applying 2, setting $z=\varphi(c,a^{\star})-\varphi(c,\pi(c))$ in Eq.
 6, and then using 1.
-Note that $\varepsilon^{\prime}_{w,T}=O(1/T)$ and $\widetilde{d}=\widetilde{o}(\sqrt{T})$ as long as the NN width $w$ is large enough (Verma et al., 2025).
+Note that $\varepsilon^{\prime}_{w,T}=O(1/T)$ and $\widetilde{d}=\widetilde{o}(\sqrt{T})$ as long as the NN width $w$ is large enough (Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2025).
 Above 2 and 3 show that the worst sub-optimality gap of the policy learned by Neural-ADB with UCB- and TS-based arm selection strategies decreases at a sub-linear rate with respect to the size of preference dataset, specifically at rate of $\tilde{O}((\tilde{d}/T)^{\frac{1}{2}})$, where $\tilde{O}$ hides the logarithmic factors and constants.
@@
 After receiving preference feedback for $T$ pairs of arms, the regret of a sequential arm selection policy is given by: $\mathfrak{R}_{T}=\sum_{t=1}^{T}\left[f(\varphi(c_{t},a_{t}^{\star}))-\left({f(\varphi(c_{t},a_{t,1}))+f(\varphi(a_{t,2}))}\right)/{2}\right],$ where $a_{t}^{\star}=\arg\!\max_{a\in\mathcal{A}_{t}}f(\varphi(c_{t},a))$ denotes the arm that maximizes the reward function for a given context $c_{t}$.
-In many real-world applications, such as medical treatment design (Lai & Robbins, 1985; Bengs et al., 2021) and content moderation (Avadhanula et al., 2022), both actively selecting arms and minimizing regret is required.
+In many real-world applications, such as medical treatment design (Lai and Robbins, 1985; Bengs et al., 2021) and content moderation (Avadhanula et al., 2022), both actively selecting arms and minimizing regret is required.
 For instance, in personalized medical treatment, active learning is used to identify the most informative treatments to test, while cumulative regret minimization ensures the system continually adapts to deliver better patient outcomes.
+Such scenarios also arise in other fields, such as dynamic pricing and personalized education, enabling systems to make smarter decisions, reduce suboptimal choices, and optimize overall performance as they gather more valuable data.
 Since the arm selection strategies in Neural-ADB are directly adapted from UCB- and TS-based algorithms for contextual dueling bandits of Verma et al.
@@
 For completeness, we state the regret upper bounds of Neural-ADB as follows.
-Corollary 1 (UCB).
-(Verma et al., 2025, Theorem 2) Let $\lambda>\kappa_{\mu}$ and $w\geq\text{poly}(T,L,K,1/\kappa_{\mu},L_{\mu},1/\lambda_{0},1/\lambda,\log(1/\delta))$.
-Then, with a probability of at least $1-\delta$, the regret of Neural-ADB when using UCB-based arm selection strategy is upper bounded by
-$\mathfrak{R}_{T}=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}}\right).$
-Corollary 2 (TS).
-(Verma et al., 2025, Theorem 3) Let the conditions in 1. Then, with a probability of at least $1-\delta$, the regret of Neural-ADB when using TS-based arm selection strategy is upper bounded by
+Corollary 1 (Regret Upper Bound).
+(Verma et al., 2025, Theorem 2 and Theorem 3) Let $\lambda>\kappa_{\mu}$ and $w\geq\text{poly}(T,L,K,1/\kappa_{\mu},L_{\mu},1/\lambda_{0},1/\lambda,\log(1/\delta))$.
+Then, with a probability of at least $1-\delta$, the regret of Neural-ADB when using UCB- or TS-based arm selection strategy is upper bounded by
 $\mathfrak{R}_{T}=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}}\right).$
@@
 To validate our theoretical results, we empirically evaluate the performance of our algorithms on different problem instances of synthetic datasets.
-Specifically, we use two commonly used synthetic functions adopted from existing works on neural bandits (Zhou et al., 2020; Zhang et al., 2021; Dai et al., 2023; Verma et al., 2025): $f(x)=10(x^{\top}\theta)^{2}$ (Square), $f(x)=2\sin(x^{\top}\theta)$ (Sine), and $f(x)=\cos(3x^{\top}\theta)$ (Cosine).
+Specifically, we use two commonly used synthetic functions adopted from existing works on neural bandits (Zhou et al., 2020; Zhang et al., 2021; Dai et al., 2023; Verma et al., 2025): $f(x)=10(x^{\top}\theta)^{2}$ (Square) and $f(x)=2\sin(x^{\top}\theta)$ (Sine).
 All experiments are repeated 10 times, and we report the average worst suboptimality gap with 95% confidence intervals (depicted as vertical lines on each curve).
-Additional results are also provided in Appendix A due to space constraints.
-Synthetic dataset.
-Let $x_{t,a}$ be the context-arm feature vector for context $c_{t}$ and an arm $a$.
-For all $t\geq 1$, $x_{t,a}$ is sampled uniformly at random from $(-1,1)^{d}$.
-We keep the number of arms constant across all rounds, denoted by $K$.
-In our experiments, the binary preference feedback indicating whether $x_{t,1}$ preferred over $x_{t,2}$ (representing human preference feedback) is sampled from a Bernoulli distribution with parameter $\mu\left(f(x_{t,1})-f(x_{t,2})\right)$).
-Reward function estimation.
-We use a neural network with $2$ hidden layers with width $50$ to estimate the latent reward function, $\lambda=1.0$, $\delta=0.05$, $d=20$, $K=10$, and fixed value of $\nu_{T}=\nu=1.0$ in all our experiments (unless we specifically indicate $d$ and $K$).
-Note that we did not perform any hyperparameter search for Neural-ADB.
 (a) Suboptimality Gap
@@
 (f) Average Regret
-(g) Suboptimality Gap
-(h) MAE
-(i) Average Regret
-Figure 1: Performance comparison under different synthetic functions, i.e., Square function (first row), Sine function (second row), and Cosine function (third row).
+Figure 1: Performance comparison of Neural-ADB against different active dueling bandit algorithms on synthetic functions: Square function (top row) and Sine function (bottom row).
+(a) Suboptimality Gap
+(b) MAE
+(c) Average Regret
+(d) Suboptimality Gap
+(e) MAE
+(f) Average Regret
+Figure 2: Performance of Neural-ADB (UCB) on the Square function, evaluated across varying input dimensions (top row) and numbers of arms (bottom row).
+(a) Sub-Optimality Gap
+(b) MAE
+(c) Average Regret
+(d) Sub-Optimality Gap
+(e) MAE
+(f) Average Regret
+Figure 3: Performance of Neural-ADB (TS) on the Square function, evaluated across varying input dimensions (top row) and numbers of arms (bottom row).
+Synthetic dataset.
+We generate sample features for each context-arm pair in a $d$-dimensional space.
+Let $x_{t,a}$ be the context-arm feature vector for context $c_{t}$ and an arm $a$.
+For all $t\geq 1$, $x_{t,a}$ is sampled uniformly at random from $(-1,1)^{d}$.
+We keep the number of arms constant across all rounds, denoted by $K$.
+In our experiments, the binary preference feedback indicating whether $x_{t,1}$ preferred over $x_{t,2}$ (representing human preference feedback) is sampled from a Bernoulli distribution with parameter $\mu\left(f(x_{t,1})-f(x_{t,2})\right)$), where $f$ is either a Square or Sine function..
+Reward function estimation.
+We use a neural network with $2$ hidden layers with width $50$ to estimate the latent reward function, $\lambda=1.0$, $\delta=0.05$, $d=20$, $K=10$, $T=1000$, and fixed value of $\nu_{T}=\nu=1.0$ in all our experiments (unless we specifically indicate $d$ and $K$).
+Note that we did not perform any hyperparameter search for Neural-ADB, whose performance can be further improved by doing the hyperparameter search.
 Comparison with baselines.
 We compare the worst suboptimality gap (defined in Eq.
-1), MAE (average suboptimality gap), and average regret (defined in Section 3.4) against the different baselines of active contextual dueling bandits to evaluate the performance of UCB- and TS-variant of Neural-ADB.
-We use three baselines: Random, AE-Borda (Mehta et al., 2023), APO (Das et al., 2024), AE-DPO (Mehta et al., 2023) and the neural variants of AE-Borda and AE-DPO in which we use a Neural network to estimate the latent reward function.
+1), MAE (average suboptimality gap, i.e., $\sum_{t=1}^{T}\left[\max_{a\in\mathcal{A}}f(\varphi(c_{t},a))-f\left(\varphi(c_{t},\pi(c))\right)\right]/T$), and average regret (defined in Section 3.4) against the different baselines of active contextual dueling bandits to evaluate the performance of UCB- and TS-variant of Neural-ADB.
+We use three baselines: Random, AE-Borda (Mehta et al., 2023), AE-DPO (Mehta et al., 2023), APO (Das et al., 2024), and the neural variants of AE-Borda and APO in which we use a neural network to estimate the latent reward function.
 They are named AE-Borda (NN) and Neural-ADB (APO) respectively.
 Experimental results in Fig.
-1 show that our approach – Neural-ADB (UCB) – outperforms other baselines in almost all synthetic functions (i.e., square function and sine function) in terms of the suboptimality gap.
-Moreover, our approach outperforms other baselines on all synthetic functions in MAE and average regret.
+1 show that our algorithm, Neural-ADB (UCB), outperforms other baselines in almost all synthetic functions (i.e., square function and sine function) in terms of the suboptimality gap.
+Moreover, both UCB- and TS-variants of Neural-ADB also outperform other baselines on all synthetic functions in MAE and average regret.
 We have included more comparisons of our approach with other baselines in other settings (e.g., different $d$ or $K$) in Appendix B.
+[…]
 5 Related Work
+In the following, we briefly review the relevant work, especially in neural contextual bandits and dueling bandits, to our problem setting.
+Neural Contextual Bandits.
+To model complex and non-linear reward functions, neural contextual bandits (Riquelme et al., 2018; Zhou et al., 2020; Zhang et al., 2021; Xu et al., 2022; Bae and Lee, 2025; Verma et al., 2025) use deep neural networks for reward function estimation.
+(Riquelme et al., 2018) employ multi-layer neural networks to learn arm embeddings and then use Thompson Sampling at the final layer for exploration.
+Zhou et al.
+(2020) propose the first neural contextual bandit algorithm with sub-linear regret guarantees, using a UCB exploration strategy.
+Building on this, Zhang et al.
+(2021) propose an algorithm with a TS exploration strategy.
+Ban et al.
+(2022) introduces an adaptive exploration strategy incorporating an auxiliary neural network to estimate the potential gain of the exploitation neural network, diverging from traditional UCB and TS exploration strategies.
+To reduce the computational overhead of using gradient-based features, Xu et al.
+(2022) only perform UCB-based exploration on the final layer of the neural network.
+More recent works (Bae and Lee, 2025; Verma et al., 2025) extend these techniques to handle neural contextual bandit settings with binary feedback (i.e., neural logistic bandits).
 Finite-Armed Dueling Bandits.
 Learning from preference feedback has been extensively studied in the bandit literature.
-In the finite-armed dueling bandits setting, the learner aims to find the best arm while only observing preference feedback for two selected arms (Yue & Joachims, 2009; 2011; Yue et al., 2012).
-To determine the best arm in dueling bandits, different criteria, such as the Borda winner, Condorcet winner, Copeland winner, or von Neumann winner, have been used while focusing on minimizing regret using only pairwise preference feedback (Zoghi et al., 2014b; Ailon et al., 2014; Zoghi et al., 2014a; Komiyama et al., 2015; Gajane et al., 2015; Saha & Gopalan, 2018; 2019a; 2019b; Verma et al., 2019; 2020a; 2020b; Zhu et al., 2023).
+In the finite-armed dueling bandits setting, the learner aims to find the best arm while only observing preference feedback for two selected arms (Yue and Joachims, 2009, 2011; Yue et al., 2012).
+To determine the best arm in dueling bandits, different criteria, such as the Borda winner, Condorcet winner, Copeland winner, or von Neumann winner, have been used while focusing on minimizing regret using only pairwise preference feedback (Ailon et al., 2014; Zoghi et al., 2014b, a; Gajane et al., 2015; Komiyama et al., 2015; Saha and Gopalan, 2018, 2019a, 2019b; Verma et al., 2019, 2020a, 2020b; Zhu et al., 2023).
 For a comprehensive overview of algorithms for various dueling bandits settings, we refer readers to the survey by Bengs et al.
@@
 Since the number of context-arm pairs can be potentially large or even infinite, the mean latent reward of each context-arm is assumed to be parameterized by an unknown function of its features.
-Common assumptions include linear reward models (Saha, 2021; Saha & Krishnamurthy, 2022; Bengs et al., 2022; Di et al., 2023; Li et al., 2024) and non-linear models (Verma et al., 2025).
-For our setting, we adopt the neural contextual dueling bandit algorithms proposed by Verma et al.
-(2025) to construct confidence ellipsoids for the latent non-linear reward function.
+Common assumptions include linear reward models (Saha, 2021; Bengs et al., 2022; Di et al., 2023; Saha and Krishnamurthy, 2022; Li et al., 2024) and non-linear models (Verma et al., 2025).
+For our setting, we adopt the neural contextual dueling bandit algorithms proposed in (Verma et al., 2025) to construct confidence ellipsoids for the latent non-linear reward function.
 Note that Neural-ADB can incorporate alternative confidence ellipsoids by appropriately modifying 2.
-Furthermore, our work addresses an active learning problem and analyzes the convergence rate of the worst sub-optimality gap, whereas Verma et al.
-(2025) focus on a regret minimization setting and derive upper bounds on cumulative regret.
+Furthermore, our work addresses an active learning problem and analyzes the convergence rate of the worst sub-optimality gap, whereas (Verma et al., 2025) focus on a regret minimization setting and derive upper bounds on cumulative regret.
 Active contextual dueling bandits.
 The work most closely related to ours is active contextual dueling bandit (Mehta et al., 2023; Das et al., 2024), which takes a principled approach to actively collecting preference datasets.
-However, two key differences exist between our work and existing research: the non-linear reward function and the arm-selection strategy.
+However, two key differences exist between our work and existing research: the non-linear reward function and the arm selection strategy.
 Existing studies typically assume a linear reward function, which may not be suitable for many real-world applications.
 Our work addresses this gap by extending the existing framework to incorporate non-linear reward functions in contextual dueling bandits.
-Additionally, existing approaches use different methods for selecting the pair of arms, leading to distinct arm-selection strategies compared to ours.
-As a result of these differences in both the arm-selection strategy and the non-linear reward function (which we estimate using a neural network), our analysis diverges significantly from that of prior work.
+Additionally, existing approaches use different methods for selecting the pair of arms, leading to distinct arm selection strategies compared to ours.
+As a result of these differences in both the arm selection strategy and the non-linear reward function (which we estimate using a neural network), our analysis diverges significantly from that of prior work.
 6 Conclusion
@@
 Finally, our experimental results further validate these theoretical findings.
-An interesting direction for future work is applying Neural-ADB to real-life applications such as LLM alignment (Mehta et al., 2023; Das et al., 2024).
+An interesting direction for future work is applying Neural-ADB to real-life applications such as LLM alignment.
 From a theoretical perspective, exploring the non-stationary setting presents another promising future direction.
@@
   1. 1.
-    $\lambda_{\max}(V_{T})\leq\lambda+\sum_{s=1}^{T}\lambda_{\max}\left(z_{s}z_{s}^{\top}\right)$.
+    Let $\delta\in(0,1)$, $\forall s\leq T:\ \left\|V_{s}-V_{s-1}\right\|^{2}\leq C_{s}$, where $\left\|A\right\|$ denotes the operator norm.
+Then, using Theorem 7.1 and Corollary 7.2 of Tropp (2012), with probability at least $1-\delta$,
+    $\mathbb{P}\left\{\lambda_{\max}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)\geq\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\frac{d}{\delta}\right)}\right\}\leq\delta.$
   2. 2.
-    $\lambda_{\min}(V_{T})\geq\lambda+\sum_{s=1}^{T}\lambda_{\min}\left(z_{s}z_{s}^{\top}\right)$.
+    $\lambda_{\max}(V_{T})=-{\lambda_{\min}(-V_{T})}$.
   3. 3.
-    $\lambda_{\max}\left(z_{s}z_{s}^{\top}\right)\leq\left\|z_{s}\right\|_{2}^{2}\ \implies\ \lambda_{\max}(V_{T})\leq\lambda+\sum_{s=1}^{T}\left\|z_{s}\right\|_{2}^{2}$.
-  4. 4.
-    $\lambda_{\max}(V_{T}^{-1})={1}/{\lambda_{\min}(V_{T})}$.
-  5. 5.
-    $\lambda_{\max}(V_{T})=-{\lambda_{\min}(-V_{T})}$.
-  6. 6.
     Let $\lambda_{i}(V)$ be the $i$-th eigenvalue of matrix $V$. If $W$ is any Hermitian matrix, then, from Weyl’s inequality:
@@
     $2.\ \lambda_{i}(V_{T})-\lambda_{\max}(W)\leq\lambda_{i}(V_{T}-W)\leq\lambda_{i}(V_{T})-\lambda_{\min}(W).$
-  7. 7.
-    For any vector $z\in\mathbb{R}^{d}$, the Mahalanobis norm of $z$ with respect to $V_{T}$ is $\left\|z\right\|_{V_{T}}=\sqrt{z^{\top}V_{T}z}\leq\left\|z\right\|_{2}\sqrt{\lambda_{\max}(V_{T})}$.
-  8. 8.
-    For $a>0:\ \left\|az\right\|_{V_{T}}=a\left\|z\right\|_{V_{T}}$ and $\lambda_{i}(aV_{T})=a\lambda_{i}(V_{T})$.
-  9. 9.
-    Let $z_{\max}=\arg\!\max_{z\in\mathbb{R}^{d}}\left\|z\right\|_{V_{T}}$. Then, $z_{\max}$ is the eigenvector of $V_{T}$ corresponding to its largest eigenvalue.
-  10. 10.
+  4. 4.
     Let $\forall z\in\mathbb{R}^{d}\colon\ \left\|z\right\|_{2}\leq L$.
 Then, $\max_{z\in\mathbb{R}^{d}}\left\|z\right\|_{V_{T}^{-1}}\leq\left\|z\right\|_{2}\sqrt{\lambda_{\max}(V_{T}^{-1})}\leq L/\sqrt{\lambda_{\min}(V_{T})}$.
-  11. 11.
-    Let $\delta\in(0,1)$, $\forall s\leq T:\ \left\|V_{s}-V_{s-1}\right\|^{2}\leq C_{s}$, where $\left\|A\right\|$ denotes the operator norm.
-Then, using Theorem 7.1 and Corollary 7.2 of Tropp (2012), with probability at least $1-\delta$,
-    $\mathbb{P}\left\{\lambda_{\max}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)\geq\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\frac{d}{\delta}\right)}\right\}\leq\delta.$
+  5. 5.
+    For $a>0:\ \left\|az\right\|_{V_{T}}=a\left\|z\right\|_{V_{T}}$ and $\lambda_{i}(aV_{T})=a\lambda_{i}(V_{T})$.
 A.1 Proof of 1
@@
 Proof.
-Using Property 11 from 1 with $Y_{T}-\mathbb{E}\left[Y_{T}\right]=\mathbb{E}\left[V_{T}\right]-V_{T}$, we have
+Using Property 1 in 1 with $Y_{T}-\mathbb{E}\left[Y_{T}\right]=\mathbb{E}\left[V_{T}\right]-V_{T}$, we have
 $\mathbb{P}\left\{\lambda_{\max}\left(\mathbb{E}\left[V_{T}\right]-V_{T}\right)\geq\tau\right\}\leq d\exp\left(\frac{-\tau^{2}}{8\sum_{s=1}^{T}C_{s}}\right)$
@@
 $\implies$ $\mathbb{P}\left\{\lambda_{\min}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)\leq-\tau\right\}\leq d\exp\left(\frac{-\tau^{2}}{8\sum_{s=1}^{T}C_{s}}\right).$
-Using upper bound on $\lambda_{\min}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)$ from Property 6 in 1, we get
+Using upper bound on $\lambda_{\min}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)$ from Property 3 in 1, we get
 $\implies$ $\mathbb{P}\left\{\lambda_{\min}(V_{T})-\lambda_{\min}(\mathbb{E}\left[V_{T}\right])\leq-\tau\right\}\leq d\exp\left(\frac{-\tau^{2}}{8\sum_{s=1}^{T}C_{s}}\right)$
@@
 Therefore, with probability at least $1-\delta$, $\lambda_{\min}(V_{T})\geq T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\nicefrac{{d}}{{\delta}}\right)}$.
-Using Property 10 from 1, we now use to prove our key result as follows:
+Using Property 4 in 1, we now use to prove our key result as follows:
 $\max_{z\in\mathcal{Z}}\left\|z\right\|_{V_{T}^{-1}}$ $\leq L/\sqrt{\lambda_{\min}(V_{T})}$
@@
 $\implies\max_{z\in\mathcal{Z}}\left\|z\right\|_{V_{T}^{-1}}$ $\leq=L/G_{T}.\qed$
-A.2 Proof of 1
+A.2 Proof of 1 and 2
 Our next results gives an upper bound of worst sub-optimality gap in terms of the upper bound of estimation error in the reward difference between any triple of context and two arms. See 1
@@
 ∎
+See 2
+Proof.
+Recall that we are using the arm-selection strategies proposed in (Verma et al., 2025).
+Since their confidence bounds hold for any adapted sequence of contexts, the proof of the first part follows directly from Theorem 1 in (Verma et al., 2025), while the second part follows from Lemma 10 together with Eq.
+(27) of (Verma et al., 2025).
+∎
+Remark 1.
+We adopt the arm selection strategies from the existing neural dueling bandit algorithms in (Verma et al., 2025), which assume $\tilde{d}=o(T)$.
+In some cases, $\tilde{d}=\Omega(T)$ (Ban et al., 2022; Deb et al., 2024), which may result in a constant convergence rate.
+However, our objective is to demonstrate the use of the neural network for estimating non-linear reward functions in active contextual dueling bandits.
+Since neural dueling bandit algorithms primarily influence the arm selection strategy, we can incorporate any variants of these algorithms by making appropriate modifications to 2.
 A.3 Proof of 2 and 3
@@
 […]
-[…]
+$=\nu_{T}\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\max_{c\in\mathcal{C}}\left(\left\|\varphi(c,a^{\star})-\varphi(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
 […]
@@
 $\Delta_{T}^{\pi}$ $\leq\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sigma_{T}(c,a^{\star},\pi(c))\right)+2\varepsilon^{\prime}_{w,T}$
-$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}}}\left\|\frac{\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))}{\sqrt{w}}\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
-$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\left\|\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
-$=\nu_{T}\log\left(KT^{2}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\max_{c\in\mathcal{C}}\left(\left\|\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
+$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}}}\left\|\frac{\varphi(c,a^{\star})-\varphi(c,\pi(c))}{\sqrt{w}}\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
+$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\left\|\varphi(c,a^{\star})-\varphi(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
+$=\nu_{T}\log\left(KT^{2}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\max_{c\in\mathcal{C}}\left(\left\|\varphi(c,a^{\star})-\varphi(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
 $\leq\nu_{T}\log\left(KT^{2}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\left(\frac{L}{\sqrt{T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\frac{d}{\delta}\right)}}}\right)+2\varepsilon^{\prime}_{w,T}$
 $\leq\tilde{O}\left(\sqrt{\frac{\tilde{d}}{T}}\right).$
-Appendix B Experimental Details and Additional Results
+Appendix B Additional Experimental Details and Results
 B.1 Experimental Details
@@
 Practical considerations.
-Based on the neural tangent kernel (NTK) theory (Jacot et al., 2018), the initial gradient $g(x;\theta_{0})$ can be used as the original feature vector $x$ as $g(x;\theta_{0})$ effectively represents the random Fourier features of the NTK.
+Based on the neural tangent kernel (NTK) theory (Jacot et al., 2018), the initial gradient $g(x;\theta_{0})$ can be used as serve as a surrogate for the original feature vector $x$ as $g(x;\theta_{0})$ effectively represents the random Fourier features of the NTK.
 To make our algorithm more practical, we use common practices in neural bandits (Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2025).
@@
 We retrain the neural network after every $20$ rounds for $50$ gradient steps across all experiments.
-B.2 Additional Results
-Next, we present the additional experiment results comparing the performance of our proposed algorithms across varying input dimensions and different numbers of arms.
-We then evaluate the performance of Neural-ADB (TS) under various input dimensions and numbers of arms.
+B.2 Additional Experimental Results
+Next, we present the additional experiment results comparing the performance of Neural-ADB varying input dimension $d$ (Fig. 4) and different numbers of arms $K$ (Fig. 5).
 (a) Sub-Optimality Gap
@@
 (f) Average Regret
-Figure 2: Performance comparison under different input dimensionalities, i.e., 20-dimensional input (first row), 40-dimensional input (second row).
-The arm size is 10.
-We are using Square function for these results.
+Figure 4: Performance comparison across different input dimensions $d$: $d=20$ (first row) and $d=40$ (second row). We set the number of arms to $10$ and use the Square function for all experiments.
 (a) Sub-Optimality Gap
@@
 (i) Average Regret
-Figure 3: Performance comparison under different numbers of arms.
-The first row is when the arm size is 5, and the second row is when the arm size is 10, the third row is when the arm size is 15.
-The dimensionality of input is 20.
-We are using Square function for these results.
-[…]
-(a) Suboptimality Gap
-(b) MAE
-(c) Average Regret
-(d) Suboptimality Gap
-(e) MAE
-(f) Average Regret
-Figure 4: Performance of Neural-ADB (UCB) under different input dimensions (first row) and numbers of arms (second row). We are using Square function for these results.
+Figure 5: Performance comparison across different numbers of arms $K$: $K=5$ (top row), $K=10$ (middle row), and $K=15$ (bottom row).
+We set the input dimension to $20$ and use the Square function for all experiments.
+Performance vs.
+neural network size.
+To investigate how performance varies with different neural network (NN) sizes, we used the Square and Cosine functions defined in the paper.
+We varied either the number of layers (with width = 32) or the width of the NN (with 2 layers), while keeping all other variables consistent with those in the paper.
+As shown in Fig.
+6, we observed that selecting the appropriate size of NN is crucial for the given problem.
+Using a large NN for a simple problem leads to poor performance due to high bias in the estimation, while a smaller NN may not accurately be able to estimate the complex non-linear function.
 (a) Sub-Optimality Gap
@@
 (f) Average Regret
-Figure 5: Performance of Neural-ADB (TS) under different input dimensions (first row) and numbers of arms (second row). We are using Square function for these results.
+Figure 6: We compare performance across different neural network widths (first row) and numbers of hidden layers (second row), using the Square function in all experiments.
+All other parameters are kept fixed, except that the width is set to $32$ when varying the number of layers.
+B.3 Computational Efficiency.
+To discuss the computational efficiency of Neural-ADB, we follow the approach of (Verma et al., 2025) and consider the following two key aspects: size of the neural network and then the number of contexts and arms.
+Size of the neural network.
+The primary computational cost in Neural-ADB arises from the neural network (NN) used to approximate the latent non-linear reward function.
+Given a context-arm feature vector of dimension $d$, an NN with $D$ hidden layers and $w$ neurons per layer incurs an inference cost of $O(dw+Dw^{2}+w)$ per context-arm pair.
+The total number of parameters in the NN is $p=dw+Dw^{2}+w$, and the training time per iteration is $O(\mathcal{E}\mathcal{P}Dw^{2})$, where $\mathcal{E}$ is the number of training epochs and $\mathcal{P}$ is the number of observed context-arm pairs.
+Choosing an appropriate NN size is critical, as NNs that are too small may fail to accurately approximate the underlying non-linear reward function, while excessively large NNs can result in substantial training and inference overhead.
+Number of contexts and arms.
+Let $K$ denote the number of arms and $p$ the total number of NN parameters.
+Since Neural-ADB uses NN gradients as context-arm features, the cost of computing gradients for all arms per context is $O(K^{2}dp)$, where $d$ is the dimension of the context-arm feature vector.
+The cost of computing reward estimates and confidence terms for all context-arm pairs is $O(K^{2}p)$ and $O(K^{2}p^{2})$, respectively.
+For arm selection, the first selection step requires $O(Kp+K)$, consisting of reward estimation for all arms ($O(Kp)$) and then identifying the arm with the highest estimated reward ($O(K)$).
+The second arm selection incurs a cost of $O(Kp+(K-1)p^{2})$, including reward estimation $O(Kp)$ and confidence term computation $O((K-1)p^{2})$ relative to the first selected arm.
+Thus, the total computational cost for selecting a pair of arms per context is $O(K^{2}dp+K^{2}p^{2})$.
+Since each context-arm pair is independent, gradients, reward estimates, and optimistic terms can be computed in parallel, reducing the overall cost to $O(dp+p^{2})$ for each iteration.